Theory and Applications of Categories, Vol. 24, No. 13, 2010, pp. 302–377.

THE SPAN CONSTRUCTION

ROBERT DAWSON, ROBERT PARÉ, DORETTE PRONK

Abstract. We present two generalizations of the Span construction. The first generalization gives Span of a category with all pullbacks as a (weak) double category. This double category 𝕊pan A can be viewed as the free double category on the vertical category A where every vertical arrow has both a companion and a conjoint (and these companions and conjoints are adjoint to each other). Thus defined, 𝕊pan: Cat → Doub becomes a 2-functor, which is a partial left bi-adjoint to the forgetful functor Vrt : Doub → Cat, which sends a double category to its category of vertical arrows.

The second generalization gives Span of an arbitrary category as an oplax normal double category. The universal property can again be given in terms of companions and conjoints and the presence of their composites. Moreover, 𝕊pan A is universal with this property in the sense that 𝕊pan: Cat → OplaxNDoub is left bi-adjoint to the forgetful functor which sends an oplax double category to its vertical arrow category.

1. Introduction

As remarked in [12], the 2-category Π_{2}A obtained by freely adjoining a right adjoint to each arrow of a category A, studied in [14] and [15], has zig-zag paths of arrows of A as 1-cells and equivalence classes of fences as 2-cells. Thus, it would appear that this construction can be performed in two steps: first, take spans in A (that is, one step zigzags), and then take paths of these. Moreover, each of these constructions is interesting in itself and deserving of further study. And indeed this is so. Paths and spans are also essential building blocks for other localizations such as the hammock localization [10] and categories of fractions [17].

In [12] we made a systematic study of the path construction. There we saw that it was useful, even necessary, to work in the context of double categories and even Leinster’s fc-multicategories (which we called, more suggestively, lax double categories). The more important construction presented there is ℙath_{∗}, which arises as the solution of the problem of describing the universal oplax normal morphism of double categories. It is in this construction that the equivalence relation on cells in Π_{2}A first makes its appearance, and this fact certainly clarifies part of the role of the equivalence relation.

All three authors are supported by NSERC discovery grants. The third author also thanks Macquarie University for its hospitality and financial support during a sabbatical visit.

Received by the editors 2009-12-31 and, in revised form, 2010-07-21. Transmitted by Ronald Brown. Published on 2010-07-23. 2000 Mathematics Subject Classification: 18A40, 18C20, 18D05. Key words and phrases: Double categories, Span construction, Localizations, Companions, Conjoints,

Adjoints.

\begin{matrix}c \\ ◯ \end{matrix} Robert Dawson, Robert Paré, Dorette Pronk, 2010. Permission to copy for private use granted.

To complete our construction of Π_{2}A we must study spans in A. If A is a category with pullbacks, the bicategory Span A is well-known. It was introduced by Bénabou in 1967; from there, it is an easy step to pass to the (weak) double category 𝕊pan A, first introduced in [19], and briefly considered in [13]. This double category of spans is better behaved than the bicategory when it comes to universal properties.

However, localization constructions such as Π_{2} typically do not require categories to have pullbacks. Without pullbacks we cannot compose spans, but we do know what a cell from a span into a path of spans ought to be, because the composite of a path of spans is a limit construction. Generalizations of the Span construction are among the main motivating examples of Hermida’s multicategories with several objects [21], and Leinster’s fc-multicategories [26]. It is our purpose in this paper to elevate these versions of Span from mere examples to the central object of study.

It is clear that all these Span constructions are natural, but it would be reassuring if we had a universal property to validate them, and provide a framework for further extensions and a better understanding of the relationship between the Span construction and other localization constructions. In [13] we established two universal properties of Span A, one of which did not refer to pullbacks. We will extend these properties in the present context.

In Section 2 we review the properties of Span as a bicategory construction, and give a slightly improved version of our characterization of gregarious morphisms from [13] (which were called ‘jointed morphisms’ in that paper). In Section 3 we describe the universal properties of the 2-functor 𝕊pan: Cat → Doub, defined for categories with pullbacks. We introduce the notions of gregarious double categories and Beck-Chevalley double categories. In Section 4 we extend the 𝕊pan construction to arbitrary categories and obtain the aforementioned oplax double categories.

2. The bicategory of spans

We begin by recalling the basic facts about spans. If A is a category with pullbacks, the bicategory Span A has the same objects as A, but its morphisms are spans,

\xymatrix{A & S \ar[l]_{p} \ar[r]^{q} & B}.  (1)

We will denote such a span by a pair (p, q), or by its middle object S when there is no confusion possible. The composite T ⊗_{B} S, also written as T S, of \xymatrix{A & S \ar[l]_{p} \ar[r]^{q} & B} with \xymatrix{B & T \ar[l]_{r} \ar[r]^{s} & C} is given by the pullback

[figure: T ⊗_{B} S; Pb; S; p; q; T; r; s; A; B; C]

and the identity I_{A} on an object A is given by \xymatrix{A & A \ar[l]_{1_{A}} \ar[r]^{1_{A}} & A.} Composition, being defined as it is by a universal property, can hardly be expected to be associative or unitary on the nose, but it is up to coherent isomorphism. In order to express this we must have morphisms between spans.

A 2-cell t: S → S′ is a commutative diagram

[figure: A; S; p; q; t; B; A; S′; p′; q′; B.]

Vertical composition of 2-cells is defined in the obvious way, whereas horizontal composition can be defined using the universal property of the pullbacks used in the composition of the spans. So we obtain a bicategory Span A. This construction was the main object of study in our previous paper [13]. In the current paper we will revisit this construction to discuss its functoriality properties and then we will introduce and study several extensions.

The category A embeds into Span A via a morphism of bicategories

( )_{∗}: A → Span A,

which assigns to the morphism f: A → B the span f_{∗} = (\xymatrix{A & A \ar[l]_{1_{A}} \ar[r]^{f} & B).} In saying that ( )_{∗} is a morphism of bicategories we are considering A as a bicategory with only identity 2-cells. Depending on our choice of pullbacks, a point on which we don’t want to commit ourselves, composition may only be preserved up to coherent isomorphism, (gf)_{∗} ≅ g_{∗}f_{∗}. The morphism ( )_{∗} is locally full and faithful which means that there is no 2-cell f_{∗} ⇒ f′_{∗} unless f = f′ and then there is only the identity.

There is another embedding, this time contravariant, of A into Span A,

( )^{∗}: A^{op} → Span A,

which takes a morphism f: A → B to the span f^{∗} = (\xymatrix{B & A \ar[r]^{1_{A}} \ar[l]_{f} & A).}

The relationship between the images f_{∗} and f^{∗} is one of adjointness. Span A, being a bicategory, supports the notion of adjointness, expressed in terms of unit and counit 2-cells satisfying the triangle identities as described in [20], page 137; in Span A, f_{∗} is left adjoint to f^{∗}. All of these spans are further related by the Beck-Chevalley condition: If

\xymatrix{P \ar[d]_{h} \ar[r]^{k} & C \ar[d]^{g} \\ A \ar[r]_{f} & B}

is a pullback diagram in A, then the canonical 2-cell

[figure: P; k_{∗}; ⇒; C; A; h^{∗}; f_{∗}; B; g^{∗}]

is an isomorphism.

Every span \xymatrix{A & S \ar[l]_{p} \ar[r]^{q} & B} is isomorphic to q_{∗} ⊗_{S} p^{∗} so that the arrows of Span A are generated by the arrows of A (i.e., the arrows of the form f_{∗} for f ∈ A) and their right adjoints. It is also true, although less obvious, that each 2-cell of Span A can be factored as a composition of identities, units and counits of adjunctions f_{∗} ⊣ f^{∗} and Span A is in some sense the free bicategory generated by the arrows of A and their adjoints. However, it is not so in the obvious way, because it is not equivalent to Π_{2}A of [14]. Indeed, it was the object of [13] to make this precise, but before we can state the main theorem from that paper we must recall some definitions.

2.1. Gregarious morphisms of bicategories. From the beginning, [2], it was realized that it was too much to require that morphisms of bicategories preserve composition, even up to isomorphism. There are many examples to back this up. We give just one which will be important to us later.

Let A and B be categories with pullbacks and F: A → B an arbitrary functor. This gives rise to Span(F): Span A → Span B defined on objects, spans, and 2-cells by the application of F. So F assigns to the 2-cell

[figure: A; S; p; q; t; B; A; S′; p′; q′; B,]

the 2-cell

[figure: F A; F S; F p; F q; F t; F B; F A; F S′; F p′; F q′; F B.]

Vertical composition of 2-cells is preserved by this functor, but composition of spans will not be, even up to isomorphism, unless F sends pullbacks to pullbacks. In general, the universal property of pullbacks only gives us a comparison cell φ_{T,S}: F (T ⊗_{B} S) → F T ⊗_{F B} F S. Such an F will be an oplax morphism of bicategories (the vertical dual of Bénabou’s notion of morphism of bicategories [2]).

2.2. Remark. One might object that we should not consider functors F that don’t preserve pullbacks in our study of the Span construction, and there is some truth to that. But as noted in [7], one might encounter a pullback-preserving functor U: B → A to which we might want to apply Span, and this U could have a left adjoint F which does not preserve pullbacks, a common and important situation. Moreover, if this F induces some kind of left adjoint Span A → Span B to Span(U), we should take it into consideration.

To gain further insight into these questions we will address the 2-functoriality of the Span construction at the end of this section, and further at the end of Sections 3 and 4.

The oplax morphisms Span(F): Span A → Span B are all normal. In fact, any morphism of the form Span(F) preserves identities on the nose, although this in itself is not enough to ensure normality. The identities must satisfy the unit laws for structure cells of oplax morphisms of bicategories, which they do. Normality plays a fundamental role throughout this paper. One of the features of normal morphisms is that one can easily characterize when they preserve adjunctions.

We say that an oplax morphism of bicategories Φ: A → B preserves adjunctions if for each adjunction \xymatrix{A \ar[r]|{f} & A′ \ar[l]|{u}} in A with unit η: 1_{A} → uf and counit ε: fu → 1_{A′}, there are 2-cells η̄: 1_{ΦA} → Φ(u)Φ(f) and ε̄: Φ(f)Φ(u) → 1_{ΦA′} which satisfy the triangle identities and give rise to commutative squares

\xymatrix{Φ(1_{A}) \ar[r]^{Φη} \ar[dd]_{ϕ_{A}} & Φ(uf) \ar[dd]^{ϕ_{u,f}} &  & Φ(fu) \ar[dd]_{ϕ_{f,u}} \ar[r]^{Φε} & Φ(1′_{A}) \ar[dd]^{ϕ_{A′}} \\  &  & and \\ 1_{ΦA} \ar[r]_{η̄} & Φ(u)Φ(f) &  & Φ(f)Φ(u) \ar[r]_{ε̄} & 1_{ΦA′} .}

So if Φ is normal we can express η̄ directly in terms of Φη and the structure cells. However, to define ε̄ in terms of Φε we need that Φ preserves the composite fu up to isomorphism. Proposition 2.3 below shows that in the presence of normality this condition is also sufficient to ensure that Φ preserves adjunctions.

We see from this discussion that for bicategories normality is a useful property when considering the preservation of adjunctions, but it is not necessary or sufficient to ensure this preservation. We will see in Proposition 3.8 that in the context of double categories normality is equivalent to the preservation of conjoints, a kind of adjointness between horizontal and vertical morphisms.

The oplax morphisms of the form Span(F) are more than normal, they preserve not only identities, but also certain composites, namely those of the form f_{∗} ⊗ S and those of the form S ⊗ g^{∗}, for any span S. Indeed, the pullback involved in the composite f_{∗} ⊗ S can be chosen to be

[figure: S; q; S; q; B; B]

and such pullbacks are preserved by any functor. The case of S ⊗ g^{∗} is dual. We saw in [13] that the condition that an oplax normal morphism preserve composites of the form fx for every left adjoint f is sufficient to ensure the preservation of adjunctions. However, there is the following slightly stronger result.

2.3. Proposition. Let Φ: X → Y be an oplax normal morphism of bicategories and let f ⊣ u be an adjoint pair of arrows in X. If Φ preserves the composite fu, then Φ(f) is left adjoint to Φ(u). Furthermore, in this case Φ preserves all composites of the form fx and zu.

Proof. Let ε: fu → 1_{B} and η: 1_{A} → uf be the adjunctions for f ⊣ u. To say that Φ preserves the composite fu means that the oplaxity 2-cell φ_{f,u}: Φ(fu) → Φ(f)Φ(u) is invertible. Define

\overline{ε} = (\xymatrix{Φ(f)Φ(u) \ar[r]^{φ_{f,u}^{−1}} & Φ(fu) \ar[r]^{Φ(ε)} & Φ(1_{B}) \ar[r]^{φ_{B}} & 1_{ΦB}})

and

\overline{η} = (\xymatrix{1_{ΦA} \ar[r]^{φ_{A}^{−1}} & Φ(1_{A}) \ar[r]^{Φ(η)} & Φ(uf) \ar[r]^{φ_{u,f}} & Φ(u)Φ(f)}).

Then we obtain commutative diagrams

[figure: Φ(f); ∼; Φ(f)1_{ΦA}; Φ(f)φ_{A}^{−1}; Φ(f)Φ(1_{A}); coh; φ_{f,1_{A}}^{−1}; Φ(f)Φ(η); Φ(f)Φ(uf); nat; coh; Φ(f)φ_{u,f}; Φ(f)Φ(u)Φ(f); φ_{f,u}^{−1}Φ(f); Φ(f1_{A}); coh; Φ(fη); Φ(fuf); Δ; φ_{f,uf}; φ_{fu,f}; Φ(εf); Φ(fu)Φ(f); nat; Φ(ε)Φ(f); Φ(1_{B}f); φ_{1_{B} ,f}; Φ(1_{B})Φ(f); coh; φ_{B}Φ(f); 1_{ΦB}Φ(f); ≀; Φ(f)]

and

[figure: Φ(u); ∼; 1_{ΦA}Φ(u); φ_{A}^{−1}Φ(u); Φ(1_{A})Φ(u); coh; φ_{1_{A},u}^{−1}; Φ(η)Φ(u); Φ(uf)Φ(u); nat; coh; φ_{u,f} Φ(u); Φ(u)Φ(f)Φ(u); Φ(u)φ_{f,u}^{−1}; Φ(1_{A}u); coh; Φ(ηu); Φ(ufu); Δ; φ_{uf,u}; φ_{u,fu}; Φ(uε); Φ(u)Φ(fu); nat; Φ(u)Φ(ε); Φ(u1_{B}); φ_{u,1_{B}}; Φ(u)Φ(1_{B}); coh; Φ(u)φ_{B}; Φ(u)1_{ΦB}; ≀; Φ(u)]

where the diagonal arrows come from unit isomorphisms. Note the arrows φ_{f,uf} and φ_{uf,u} in the middle which go up. This is necessary, because we do not know whether they are invertible. However, the φ_{f,1} and φ_{1,u} are invertible, since they are one-sided inverses to Φ(f)φ_{A} and φ_{A}Φ(u) respectively, which are invertible by normality. Thus, Φ(f) ⊣ Φ(u).

Next we will show that composites of the form fx are preserved (up to isomorphism). The inverse for φ_{f,x}: Φ(fx) → Φ(f)Φ(x) is the composite on the left of the diagram

[figure: Φ(f)Φ(x); ≀; Φ(f)Φ(1_{A}x); Φ(f)Φ(ηx); Φ(f)φ_{1_{A},x}; nat; Φ(f)Φ(1_{A})Φ(x); Φ(f)Φ(η)Φ(x); Φ(f)Φ(ufx); Φ(f)φ_{u,fx}; Φ(f)φ_{uf,x}; coh; Φ(f)Φ(uf)Φ(x); Φ(f)φ_{u,f} Φ(x); Φ(f)Φ(u)Φ(fx); \overline{ε}Φ(fx); Φ(f)Φ(u)φ_{f,x}; nat; Φ(f)Φ(u)Φ(f)Φ(x); \overline{ε}Φ(f)Φ(x); 1_{ΦB}Φ(fx); ≀; 1_{ΦB}φ_{f,x}; nat; 1_{ΦB}Φ(f)Φ(x); ≀; Φ(fx); φ_{f,x}; Φ(f)Φ(x)]

and Φ(f)φ_{1_{A},x} = Φ(f)φ_{A}^{−1}Φ(x) by normality and coherence of φ. Using the first triangle identity we conclude that the right side of the previous diagram, preceded by the top is id_{Φ(f)Φ(x)}.

In the other direction we have the diagram

[figure: Φ(fx); ≀; nat; φ_{f,x}; Φ(f)Φ(x); ≀; Φ(f1_{A}x); coh; Φ(fηx); nat; φ_{f,1_{A}x}; Φ(f)Φ(1_{A}x); Φ(f)Φ(ηx); Φ(fufx); Δ; Φ(εfx); φ_{fu,fx}; coh; φ_{f,ufx}; Φ(f)Φ(ufx); Φ(f)φ_{u,fx}; Φ(1_{B}fx); nat; φ_{1_{B} ,fx}; Φ(fu)Φ(fx); Φ(ε)Φ(fx); _{def \overline{ε}}; φ_{f,u}Φ(fx); Φ(f)Φ(u)Φ(fx); \overline{ε}Φ(fx); Φ(1_{B})Φ(fx); φ_{B}Φ(fx); 1_{ΦB}Φ(fx); ≀; coh; Φ(fx).]

Preservation of zu is dual.

The result of Proposition 2.3 leads to the following definition which is fundamental to all that will follow.

2.4. Definition. An oplax morphism Φ: X → Y of bicategories is called gregarious if it is normal and preserves all composites of the form fu where f is left adjoint to u. (In [13] these morphisms were called jointed.)

The following theorem, proved in Proposition 2.8 and 2.9 of [13], establishes gregariousness as a fundamental concept.

2.5. Theorem. An oplax normal morphism of bicategories is gregarious if and only if it preserves adjoints.

The importance of this notion has been observed by various authors (cf. [6], [7], [8] and [9]). For a full discussion, see [13].

2.6. The universal property of Span. As we noted above, the embedding

( )_{∗}: A → Span(A)

sends all arrows in A to left adjoints in Span(A). So in order to state the universal property of Span we will use the following property of morphisms of bicategories, which was introduced in [13].

2.7. Definition. A (strong) morphism Φ: X → Y of bicategories is sinister if Φ(f) is a left adjoint for every arrow f of X. A strong transformation t: Φ → Ψ of sinister morphisms is called sinister if for every arrow f of X, the mate of the naturality isomorphism Ψ(f)t(X) →^{∼} t(X′)Φ(f) is itself an isomorphism t(X)Φ(f)^{∗} →^{∼} Ψ(f)^{∗}t(X′). (See [13], Section 1.2.)

The main theorem of [13] now states that ( )_{∗}: A → Span A is the universal sinister morphism in the following precise sense.

2.8. Theorem. Let A be a category with pullbacks and B an arbitrary bicategory. Then

1. Composing with ( )_{∗}: A → Span A gives an equivalence of categories between the category of gregarious morphisms Span A → B with oplax transformations and the category of sinister morphisms A → B with strong transformations.

2. Under this correspondence, strong transformations F → G: Span A ⇉ B correspond to sinister transformations F ∘ ( )_{∗} → G ∘ ( )_{∗}: A ⇉ B.

3. Strong morphisms Span A → B correspond to sinister morphisms A → B satisfying the Beck-Chevalley condition.

2.9. Functoriality. We write PBCat for the category of categories with pullbacks and pullback-preserving functors and pbCat for the category of categories with pullbacks and arbitrary functors. Then Span can be considered either as a functor Span: PBCat → Bicat (where the morphisms of the codomain category are homomorphisms, also called pseudo functors or strong morphisms, of bicategories) or as a functor Span: pbCat → Bicat_{OplaxN} (where the morphisms of the codomain category are oplax normal morphisms of bicategories). Since we are interested in the preservation of adjunctions, we may want to ask the question whether either of these functors is a 2-functor.

However, this question does not make sense because neither Bicat_{OplaxN} nor Bicat can be considered as a 2-category (or bicategory) in a way that would include natural transformations with non-identity components on the objects and non-identity 2-cells for the arrows. For Bicat_{OplaxN} this is obvious as there is no way of whiskering any type of natural transformations with oplax morphisms. For Bicat with homomorphisms there is a notion of whiskering. However, there are several problems with the horizontal and vertical composition of the natural transformations: vertical composition is not necessarily associative on the nose if the codomain object of the homomorphisms is a bicategory, and there are problems with middle-four-interchange and horizontal composition as well. These issues have been discussed in detail in [24].

We will revisit the question of functoriality of the Span construction again at the end of the next section and show that with a slight modification of this construction the problem can be resolved and we will gain further insight into the existence of certain limits and colimits and their properties in the Span categories.

3. The double category of spans

There is another version of the Span construction which will be central to what follows. Although a seemingly minor modification of Span A, the difference will be crucial. We are referring to the (weak) double category 𝕊pan A.

As before, let A be a category with pullbacks. 𝕊pan A is the double category with the same objects as A. The vertical arrows are the morphisms of A, the horizontal arrows are the spans in A and the double cells are commutative diagrams

\xymatrix{A \ar[d]_{f} & S \ar[l]_{p} \ar[d]_{x} \ar[r]^{q} & B \ar[d]^{g} \\ A′ & S′ \ar[l]^{p′} \ar[r]_{q′} & B′.}

Vertical composition is as in A and horizontal composition as in Span A. Of course, horizontal composition is not strictly unitary or associative, so we only get a weak double category. This is most easily expressed by saying that the substructure consisting of objects, horizontal arrows, and (horizontally) special cells, form a bicategory. Here, special cells are those which have identity arrows as horizontal domain and codomain, so they are of the form

[figure: A; S; p; x; q; B; A; S′; p′; q′; B.]

So there are special associativity and unit cells that satisfy the usual coherence conditions. Where this doesn’t lead to confusion we will suppress the occurrence of these cells in our pasting diagrams. Note that, while horizontal composition of spans requires a choice of pullbacks and is only associative up to isomorphism once this choice has been made, the horizontal composition of cells is uniquely determined and as associative as possible.

3.1. Convention. For the rest of this paper, except for Section 4.28, we will refer to weak double categories simply as double categories as they seem to be the more central concept. We will talk of a strict double category when the bicategory substructure mentioned above is in fact a 2-category.

An important aspect of this new Span construction is that it retains the morphisms of A in the structure. And these can be used to formulate concepts and universal properties which Span A doesn’t support, as we shall see later. Shulman has also noted this important feature in [30].

Let us digress a bit to elaborate on this. It is useful to view internal categories in A as monads in Span A. However, this monad approach does not give us internal functors as a natural notion of morphisms between internal categories. Looking for guidance, we might consider a monad in Span A as a lax morphism defined on the bicategory 1 and so we might consider lax, oplax, or strong transformations between them. However, none of these give functors (nor do any of them give profunctors). In order to get internal functors the usual trick is to restrict to lax transformations whose components are maps. While this works, it is certainly ad hoc and does not generalize properly. For example, in the bicategory V -Mat of V -matrices, monads are V -categories, but maps between V -monads do not give V -functors.

Taking the double category 𝕊pan A instead, lax morphisms 1 → 𝕊pan A are still category objects in A, but now we have a good notion of vertical transformation between lax morphisms, and these are exactly the internal functors. Note in passing that lax morphisms 2→ 𝕊pan A are profunctors, and that vertical transformations between them are cells in the double category ℂat of categories, functors, and profunctors.

3.2. Oplax morphisms of double categories. We mentioned lax morphisms of double categories and we should say a word about the definitions of lax and oplax morphisms. It is spelled out in detail in [12], Section 1.9, for the oplax version (which is most useful for our work). An oplax morphism of double categories F: 𝔸 → 𝔹 assigns objects, horizontal arrows, vertical arrows, and double cells of 𝔹 to similar ones in 𝔸 respecting all domains and codomains and preserving vertical composition (of arrows and cells). Horizontal identities and composition are not preserved, but comparison special cells

[figure: F A; F 1_{A}; φ_{A}; F A; and; F A; F (f′f); φ_{f′,f}; F A′′; F A; 1_{F A}; F A; F A; F f; F A′; F f′; F A′′]

are given, satisfying the coherence conditions given in Section 1.9 of [12]. There are three conditions which are straightforward generalizations of the coherence conditions for oplax morphisms of bicategories, and one condition proper to double categories, namely naturality of ϕ_{A} in A, with respect to vertical arrows: For every vertical arrow \xymatrix{v: A \ar[r]|{•} & B} we require

[figure: F A; •; F v; F 1_{v}; F 1_{A}; F A; •; F v; F A; F 1_{A}; φ_{A}; F A; F B; F 1_{B}; φ_{B}; F B; =; F A; •; F v; 1_{F v}; 1_{F A}; F A; •; F v; F B; 1_{F B}; F B; F B; 1_{F B}; F B.]

Note that this condition is vacuous for bicategories (considered as double categories with only identity vertical arrows).

Vertical transformations t: F → G between oplax morphisms were also defined in [12], Definition 1.13. They associate to each object A of 𝔸 a vertical arrow \xymatrix{t_{A}: F A \ar[r]|{•} & GA} and to each horizontal arrow f: A → A′ of 𝔸 a cell

[figure: F A; F f; •; t_{A}; t_{f}; F A′; •; t_{A′}; GA; Gf; GA′]

satisfying four conditions, two saying that t is vertically natural and two saying that it is horizontally compatible with the identities and composition.

In this way we get a 2-category Doub_{Opl} of double categories, oplax morphisms, and vertical transformations. Because oplax morphisms are vertically functorial and vertical transformations are vertically natural we get a canonical forgetful 2-functor

Vrt : Doub_{Opl} → Cat,

which forgets everything but the vertical structure. It can be thought of as the ‘category of objects’ functor, generalizing Ob: Cat → Set.

3.3. Gregarious Double Categories. Returning to our general discussion of 𝕊pan A, apart from ( )_{∗} and ( )^{∗} there is a third, more direct, way of including A in 𝕊pan A, viz. the inclusion of A as vertical arrows. We will always consider A as contained in 𝕊pan A in this way and not use any special notation for the inclusion.

Now the question is, what is the relationship between f, f_{∗}, and f^{∗} in 𝕊pan A? Recall from [19] the following definitions.

3.4. Definition. Let 𝔻 be a double category and consider horizontal morphisms f: A → B and u: B → A and a vertical morphism \xymatrix{v: A \ar[r]|{•} & B}. We say that f and v are companions if there exist binding cells

[figure: A; ψ; A; •; v; and; A; •; v; f; χ; B; A; f; B; B; B,]

such that

[figure: A; ψ; A; •; v; A; ψ; A; •; v; f; χ; B; =; A; f; _{id_{f}}; B; A; •; v; 1_{v}; A; •; v; and; A; f; •; v; χ; B; =; A; f; B; B; A; f; B; B; B; B; B; .]

Dually, u and v are conjoints if there exist binding cells

[figure: A; •; v; α; A; and; B; u; β; A; •; v; B; u; A; B; B,]

such that

[figure: A; •; v; α; A; B; β; u; A; •; v; α; A; =; B; u; _{id_{u}}; A; A; •; v; 1_{v}; A; •; v; and; B; β; u; A; •; v; =; B; B; u; A; B; u; A,; B; B; B; B; .]

Finally, f is left adjoint to u in 𝔻 if there exist special cells

[figure: A; η; A; and; B; ε; u; A; f; B; A; f; B; u; A; B; B,]

such that

[figure: A; η; A; f; _{id_{f}}; B; A; f; _{id_{f}}; B; A; f; _{id_{f}}; B; u; ε; A; f; B; =; A; f; B; A; f; B; B]

and

[figure: B; u; _{id_{u}}; A; η; A; B; _{id_{u}}; u; A; B; u; ε; A; f; B; u; _{id_{u}}; A; =; B; u; A; B; B; u; A; ,]

i.e., f is left adjoint to u in the bicategory of horizontal arrows and special cells.

Note that companions in symmetric double categories (where the horizontal and vertical arrow categories are the same) were studied by Brown and Mosa [4] under the name ‘connections’.

The following result extends that of Proposition 1.4 in [18].

3.5. Proposition. Given f, u, and v as above, if any two of the relations in the definition hold, so does the third.

Proof. The proof is an amusing exercise in pasting diagrams which is left to the reader. We will do one case as an example. Suppose for example that f is left adjoint to u via ε and η and that u and v are conjoints via α and β. Then define χ and ψ to be

[figure: A; •; v; α; A; f; _{id_{f}}; B; A; η; A; B; u; ε; A; f; B; and; A; f; _{id_{f}}; B; β; u; A; •; v; B; B; A; f; B; B]

respectively. Then χψ is the composite

[figure: A; η; A; _{id_{1A}}; A; _{id_{f}}; f; B; A; f; _{id_{f}}; B; β; u; A; •; v; α; A; _{if_{f}}; f; B; A; f; _{id_{f}}; B; _{id_{1B}}; B; u; ε; A; f; B; A; f; B; B; B.]

The composite αβ in the middle is equal to id_{u} so the whole middle row is an identity and the big square reduces to a vertical identity cell by the first triangle identity for η and ε.

Companions are unique up to isomorphism and compose in the sense that if f and v are companions and f′ and v′ are companions and f′ and f are composable, then f′f and v′ · v are again companions. A similar statement holds for conjoints by duality and the corresponding statement for adjoints is well known.

3.6. Examples.

1. In 𝕊pan A, f and f_{∗} are companions, the binding cells being

[figure: A; A; A; f; and; A; f; A; f; f; B; A; A; f; B; B; B; B.]

Similarly, f and f^{∗} are conjoints. The fact that f_{∗} is left adjoint to f^{∗} follows by Proposition 3.5.

2. A 2-category A gives a double category of quintets, ℚ(A), where a typical cell looks

like

[figure: A; g; f; α⇓; B; h; C; k; P .]

In ℚ(A), the horizontal f and vertical f are companions, whereas a horizontal f and a vertical v are conjoints if and only if v is left adjoint to f in A.

Normal morphisms of double categories preserve companions and conjoints, but more is true.

3.7. Definition. Let F: 𝔸 → 𝔹 be an oplax (not necessarily normal) morphism of double categories.

1. We say that F preserves companions if for each companion pair f: A → B and \xymatrix{v: A \ar[r]|{•} & B} with binding cells χ and ψ in 𝔸 there exists a double cell

[figure: F A; \overline{ψ}; F A; •; F v; F A; F f; F B]

in 𝔹, such that F ψ = \overline{ψ} · φ_{A} and F f and F v are again companions with binding cells \overline{ψ} and φ_{B} · F χ.

2. We say that F preserves conjoints if for each conjoint pair f: A → B and \xymatrix{v: B \ar[r]|{•} & A} with binding cells α and β in 𝔸 there exists a double cell

[figure: F B; •; F v; \overline{α}; F B; F A; F f; F B]

in 𝔹, such that \overline{α} · φ_{B} = F α and F f and F v are again conjoints with binding cells φ_{A} · F β and \overline{α}.

3.8. Proposition. For an oplax morphism F: 𝔸 → 𝔹 of double categories the following are equivalent:

1. F is normal;

2. F preserves companions;

3. F preserves conjoints.

Proof. To see that (1) implies (2), let \xymatrix{f: A \ar[r] & B} and \xymatrix{v: A \ar[r]|{•} & B} be companions with binding cells ψ and χ as in Definition 3.4. Let \overline{ψ} = F ψ · φ_{A}^{−1}. Then obviously, F ψ = \overline{ψ} · φ_{A}. So we need to show that

[figure: F A; φ_{A}^{−1}; 1_{F A}; F A; F A; •; F v; F χ; F f; F B; F A; F ψ; F 1_{A}; F A; •; F v; and; F B; F 1_{B}; φ_{B}; F B; F A; F f; F B; F B; 1_{F B}; F B.]

satisfy the conditions to be binding cells for F f and F v. Their vertical composition is equal to φ_{B} ·F χ·F ψ ·φ_{A}^{−1} = φ_{B} ·F (χ·ψ)·φ_{A}^{−1} = φ_{B} ·F (1_{v})·φ_{A}^{−1} = 1_{F (v)} ·φ_{A} ·φ_{A}^{−1} = 1_{F (v)}. Their horizontal composition is equal to

[figure: F A; 1_{F A}; φ_{A}^{−1}; F A; F f; _{id_{F f}}; F B; F A; F 1_{A}; F ψ; F A; •; F v; F f; F χ; F B; F A; F f; _{id_{F f}}; F B; φ_{B}; F 1_{B}; F B; F A; F f; F B; 1_{F B}; F B,]

which, by coherence for F, is equal to

[figure: F A; F f; φ_{f,1A}; F B; F A; F f; F (χψ); F B; F A; F ψ; F 1_{A}; F A; •; F v; F χ; F f; F B; =; F A; F f; φ_{1B ,f}; F B; F A; F f; _{id_{F f}}; F B; φ_{B}; F 1_{B}; F B; F A; F f; _{id_{F f}}; F B; φ_{B}; F 1_{B}; F B; F A; F f; F B; 1_{F B}; F B; F A; F f; F B; F B]

and F (χψ) = F (id_{f}) = id_{F f} and (φ_{B}id_{F f}) · φ_{1_{F f}} = id_{F f} so the whole composite is id_{F f}, i.e., F f and F v are companions with binding cells ψ̃ = F ψ · φ_{A}^{−1} and χ̃ = φ_{B} · F χ. The corresponding result for conjoints ((1) implies (3)) follows by duality.

To show that (2) implies (1), suppose that F: 𝔸 → 𝔹 preserves companions. Note that for every object A, the horizontal identity arrow 1_{A} and the vertical identity arrow Id_{A} are conjoints with binding cells χ = ψ = ι_{A}. This gives rise to the existence of a double cell

[figure: F A; 1_{F A}; _{\overline{ι}_{A}}; F A; F A; F 1_{A}; F A]

such that

φ_{A} · \overline{ι}_{A} = ι_{F A},  (2)

and F 1_{A} and the vertical identity on A are companions in 𝔹 with binding cells

[figure: F A; 1_{F A}; _{\overline{ι}_{A}}; F A; and; F A; φ_{A}; F 1_{A}; F A; F A; F 1_{A}; F A; F A; 1_{F A}; F A.]

The conditions on the binding cells give us that φ_{A}\overline{ι}_{A} = id_{F 1_{A}} and φ_{A} · \overline{ι}_{A} = ι_{F A}. Since

[figure: F A; 1_{F A}; ι_{F A}; F A; 1_{F A}; φ_{A}; F A; F A; F 1_{A}; φ_{A}; F A; F A; 1_{F A}; _{\overline{ι}_{A}}; F A; F 1_{A}; φ_{A}; F A; =; F A; 1_{F A}; _{\overline{ι}_{A}}; F A; 1_{F A}; ι_{F A}; F A; =; F A; 1_{F A}; _{\overline{ι}_{A}}; F A; F A; F 1_{A}; F A; 1_{F A}; F A; F A; F 1_{A}; F A; 1_{F A}; F A; F A; F 1_{A}; F A ,]

i.e., φ_{A}\overline{ι}_{A} = \overline{ι}_{A} · φ_{A}, we derive that \overline{ι}_{A} · φ_{A} = id_{F 1_{A}} and φ_{A} · \overline{ι}_{A} = ι_{F A}. So \overline{ι}_{A} = φ_{A}^{−1}, and we conclude that F is normal. The corresponding result for conjoints, i.e., that (3) implies (1), follows again by duality.

3.9. Proposition. An oplax normal morphism of double categories preserves composites of the form v_{∗}x, where v_{∗} is the companion of a vertical arrow v.

Proof. We need to show that the double cell

[figure: F (v_{∗}x); φ_{v∗,x}; F x; F v_{∗}]

is vertically invertible. We claim that the inverse is the pasting of the following diagram:

[figure: F (ψ_{v}id_{x}); F x; •; F v; F v_{∗}; F χ_{v}; θ_{v_{∗},x} :=; F (v_{∗}x); _{id_{F (v∗x)}}; φ_{B}; F 1_{B}; F (v_{∗}x); _{id_{F (v∗x)}}; 1_{F B}; F (v_{∗}x); .]

The following calculation shows that θ_{v_{∗},x} · φ_{v_{∗},x} = id_{F (v_{∗}x)}. By naturality of φ and the horizontal binding cell equality for companion pairs,

[figure: F (v_{∗}x); φ_{v∗,x}; F (v_{∗}x); F (χ_{v}ψ_{v}id_{x}); F (v_{∗}x); F (id_{v∗x}); F x; F (ψ_{v}id_{x}); •; F v; F (χ_{v}); F v_{∗}; F (1_{B}v_{∗}x); φ_{1B ,v∗x}; F (1_{B}v_{∗}x); φ_{1B ,v∗x}; F (v_{∗}x); _{id_{F (v∗x)}}; F 1_{B}; φ_{B}; =; _{id_{F (v∗x)}}; F (v_{∗}x); φ_{B}; F 1_{B}; =; _{id_{F (v∗x)}}; F (v_{∗}x); φ_{B}; F 1_{B}; F (v_{∗}x); _{id_{F (v∗x)}}; 1_{F B}; F (v_{∗}x); _{id_{F (v∗x)}}; 1_{F B}; F (v_{∗}x); _{id_{F (v∗x)}}; 1_{F B}; F (v_{∗}x); F (v_{∗}x); F (v_{∗}x)]

which is equal to id_{F (v_{∗}x)} by a unit law for φ.

In order to show that φ_{v_{∗},x} · θ_{v_{∗},x} = id_{F v_{∗}}id_{F x}, we first make a couple of observations about normal oplax morphisms of double categories. Let h: A → B be any horizontal arrow in 𝔸. Since φ_{A} and φ_{B} are vertically invertible in 𝔹 and

we derive that

[figure: F h; _{id_{F h}}; φ_{B}; F 1_{B}; =; F h; φ_{1_{B} ,h}^{−1}; F 1_{B}; and; F 1_{A}; φ_{A}; _{id_{F h}}; F h; =; F 1_{A}; φ_{h,1_{A}}^{−1}; F h; F h; 1_{F B}; F (1_{B}h); 1_{F A}; F h; F (h1_{A}); .] (3)

[figure: F (1_{B}h); φ_{1B ,h}; F h; _{id_{F h}}; F (h1_{A}); φ_{h,1A}; F h; _{id_{F h}}; φ_{B}; F 1_{B}; =; =; φ_{A}; F 1_{A}; F h; _{id_{F h}}; F h; 1_{F B}; F h; 1_{F A}; F h]

We can apply this to the binding cells of a companion pair to obtain the following. By naturality of φ we have

[figure: F (v_{∗}1_{A}); φ_{v∗,1_{A}}; F (v_{∗}1_{A}); F (χ_{v}ψ_{v}); F (1_{A}); F ψ_{v}; •; F v; F v_{∗}; F χ_{v}; =; φ_{1_{B} ,v∗}; F (1_{B}v_{∗}); F v_{∗}; F 1_{B}; F v_{∗}; F 1_{B}]

and by vertical composition with inverses we obtain

[figure: F 1_{A}; F ψ_{v}; •; F v; F χ_{v}; F v_{∗}; =; φ_{A}; F 1_{A}; F v_{∗}; _{id_{F v∗}}; =; φ_{A}; F 1_{A}; F v_{∗}; _{id_{F v∗}}; _{id_{F v∗}}; F v_{∗}; φ_{B}; F 1_{B}; 1_{F A}; _{id_{F v∗}}; F v_{∗}; 1_{F A}; F v_{∗}; F v_{∗}; 1_{F B}; F v_{∗}; F (χ_{v}ψ_{v}); F v_{∗}; _{id_{F v∗}}; F v_{∗}; 1_{F B}; .] (4)

We will now use this to calculate the composition φ_{v_{∗},x} · θ_{v_{∗},x}:

[figure: F (ψ_{v}id_{x}); F x; •; F v; F χ_{v}; F v_{∗}; =; F x; F (ψ_{v}id_{x}); •; F v; F χ_{v}; F v_{∗}; by associativity; _{id_{F (v∗x)}}; F (v_{∗}x); φ_{B}; F 1_{B}; F (v_{∗}x); φ_{v∗,x}; F 1_{B}; φ_{B}; F (v_{∗}x); _{id_{F (v∗x)}}; 1_{F B}; F x; F v_{∗}; 1_{F B}; F (v_{∗}x); φ_{v∗,x}; F x; F v_{∗}; =; F (1_{A}x); φ_{1_{A},x}; F v_{∗}; _{id_{F v∗}}; by naturality of φ; F x; _{id_{F x}}; F 1_{A}; F ψ; •; F v; F χ_{v}; F v_{∗}; F x; _{id_{F x}}; F v_{∗}; _{id_{F v∗}}; F 1_{B}; φ_{B}; F x; F v_{∗}; 1_{F B}]

[figure: =; F (1_{A}x); φ_{1_{A},x}; F v_{∗}; _{id_{F v∗}}; by (4); F x; _{id_{F x}}; φ_{A}; F 1_{A}; F v_{∗}; _{id_{F v∗}}; F x; F v_{∗}; 1_{F B}]

and this is equal to id_{F v_{∗}}id_{F x} by (3).

3.10. Remark. Note that we can apply Proposition 2.3 to the horizontal bicategory of a double category to show that an oplax normal morphism preserves composites of the form v_{∗}x if and only if it preserves composites of the form zv^{∗} if and only if it preserves composites of the form v_{∗}v^{∗}. So we conclude that oplax normal morphisms preserve all these composites.

3.11. Definition. A double category is gregarious if every vertical arrow has a companion and a conjoint. We write GregDoub_{OplaxN} for the 2-category of gregarious double categories, oplax normal morphisms and vertical transformations.

Gregarious double categories were introduced and studied as framed bicategories by Shulman in [30].

3.12. Examples.

1. The double category of commuting squares in a groupoid is gregarious.

2. If A is a 2-category in which every arrow has a right adjoint, then the double category ℚ (A) of quintets is gregarious.

3. For any category A with pullbacks, the double category 𝕊pan A is gregarious.

3.13. The universal property of 𝕊pan. Let 𝔸 be a gregarious double category with A_{0} as category of objects and vertical arrows and A_{1} as category of horizontal arrows and cells (with vertical composition of cells). Any choice of companions for the vertical arrows in A gives rise to a functor Ξ: A2_{0} → A_{1}, defined by:

Ξ( [figure: •; x] ) := [figure: x_{∗}]  (5)

Ξ ( [figure: •; x; •; u; •; v; •; y]) := [figure: •; u; •; u; x_{∗}; •; x; χ_{x}; ψ_{y}; •; y; •; v; •; v; y_{∗}; .]  (6)

Dually, a choice of conjoints gives rise to a functor Z: A2_{0} → A_{1}, defined by:

Z( [figure: •; x] ) := [figure: x^{∗}]  (7)

Z ( [figure: •; x; •; u; •; v; •; y]) := [figure: x^{∗}; β_{x}; •; x; •; u; •; u; •; v; •; v; •; y; α_{y}; y^{∗}; .]  (8)

As we will use the cells from these definitions frequently, we will use the following abbreviations:

ξ(x, u, v, y) = Ξ ( [figure: •; x; •; u; •; v; •; y]) and ζ(x, u, v, y) = Z ( [figure: •; x; •; u; •; v; •; y])  (9)

There are corresponding folding operations on the cells of the double category that send cells to special horizontal cells. By abuse of notation we will write for a cell

[figure: f; •; u; γ; •; v; g; ,]

ξ(γ) = [figure: ψ_{u}; f; •; u; γ; •; v; χ_{v}; v_{∗}; u_{∗}; g] and ζ(γ) = [figure: u^{∗}; β_{u}; •; u; γ; f; •; v; α_{v}; g; v^{∗}; .]

Note that ζ(ξ(x, u, v, y)) = ξ(ζ(u, x, y, v)). This cell will play an important role in our study of a Beck-Chevalley type condition for double categories, so we will give it a special name:

[figure: u^{∗}; β_{u}; •; u; •; x; χ_{x}; x_{∗}; Υ(x, u, v, y); :=; ψ_{y}; •; y; •; v; α_{v}; y_{∗}; v^{∗}]  (10)

3.14. Lemma. For any commutative cube of vertical arrows

[figure: a; x_{1}; u_{1}; b; y_{1}; v_{1}; c; d; x_{2}; u_{2}; y_{2}; v_{2}]

the cells defined in (9) and (10) satisfy

(ζ(y_{1}, b, d, y_{2})ξ(v_{1}, c, d, v_{2})) · Υ(u_{1}, x_{1}, y_{1}, v_{1}) =
Υ(u_{2}, x_{2}, y_{2}, v_{2}) · (ξ(u_{1}, a, b, u_{2})ζ(x_{1}, a, c, x_{2})).

Proof. Straightforward calculation using the cancelation properties of the companion and conjoint binding cells.

We can now state a variant of part 1 of Theorem 2.8 for the double category 𝕊pan A.

3.15. Theorem. Let A be a category with pullbacks and 𝔹 be a gregarious double category. Then composing with the inclusion A ↪ 𝕊pan A gives an equivalence of categories between the category of oplax normal morphisms 𝕊pan A → 𝔹 with vertical transformations, and the category of functors A → Vrt (𝔹) with natural transformations,

GregDoub_{OplaxN}(𝕊pan A, 𝔹) ≃ Cat (A, Vrt (𝔹)).

Proof. For every vertical arrow v in 𝔹, choose a companion v_{∗} with cells χ_{v} and ψ_{v} and a conjoint v^{∗} with cells α_{v} and β_{v}. For the vertical identity arrows, choose the companions and conjoints to be the corresponding horizontal identity arrows with the ι cells as binding cells. Note that for composable vertical morphisms v_{1} and v_{2}, we don’t get necessarily that (v_{2} · v_{1})_{∗} = (v_{2})_{∗}(v_{1})_{∗} or (v_{2} · v_{1})^{∗} = (v_{1})^{∗}(v_{2})^{∗}, but there are canonical vertical isomorphisms τ_{v_{1},v_{2}} and σ_{v_{1},v_{2}} defined by

[figure: (v_{2}·v_{1})^{∗}; σ_{v_{1},v_{2}}; (v_{2}·v_{1})^{∗}; β_{v_{2}·v_{1}}; v_{1}; v_{1}; α_{v_{1}}; :=; v_{2}; α_{v_{2}}; (v_{1})^{∗}; (v_{2})^{∗}; (v_{1})^{∗}; (v_{2})^{∗}; (v_{1})^{∗}]  (11)

and

[figure: (v_{2}·v_{1})_{∗}; τ_{v_{1},v_{2}}; ψ_{v_{1}}; v_{1}; v_{1}; (v_{2}·v_{1})_{∗}; χ_{v_{2}·v_{1}}; :=; (v_{1})_{∗}; ψ_{v_{2}}; v_{2}; (v_{1})_{∗}; (v_{2})_{∗}; (v_{1})_{∗}; (v_{2})_{∗}; .]  (12)

With these choices of companions and conjoints, a functor F: A → Vrt (𝔹) can be lifted to an oplax normal morphism F̃: 𝕊pan A → 𝔹 defined as follows. On objects and vertical arrows, F̃ = F, and on horizontal arrows, F̃(\xymatrix{A & S \ar[l]_{q} \ar[r]^{p} & B}) = (F p)_{∗}(F q)^{∗}. Further, F̃ sends a cell

\xymatrix{A_{1} \ar[d]_{v_{1}} & S_{1} \ar[l]_{q_{1}} \ar[r]^{p_{1}} \ar[d]_{w} & B_{1} \ar[d]^{v_{2}} \\ A_{2} & S_{2} \ar[l]^{q_{2}} \ar[r]_{p_{2}} & B_{2}}

in 𝕊pan A to the following horizontal composition of cells in 𝔹 (cf. (9)),

[figure: •; F v_{1}; (F q_{1})^{∗}; ζ(F q_{1},F w,F v_{1},F q_{2}); F w; •; (F p_{1})_{∗}; ξ(F p_{1},F w,F v_{2},F p_{2}); •; F v_{2}; (F q_{2})^{∗}; (F p_{2})_{∗}; .]

Note that by our choice of companions and conjoints for the vertical identity arrows, F̃ preserves horizontal identities on the nose. To obtain the comparison cells for horizontal composites, consider a composable pair of horizontal arrows

\xymatrix{A & S \ar[l]_{q} \ar[r]^{p} & B & S′ \ar[l]_{q′} \ar[r]^{p′} & C}

in 𝕊pan A, with pullback

\xymatrix{S ×_{B} S′ \ar[d]_{\overline{q′}} \ar[r]^{\overline{p}} & S′ \ar[d]^{q′} \\ S \ar[r]_{p} & B}

in A. The comparison cell φ_{S,S′} is defined as the composite

[figure: F (q\overline{q′})^{∗}; σ_{F \overline{q′},F q}; F (p′\overline{p})_{∗}; τ_{F \overline{p},F p′}; (F q)^{∗}; Υ(F \overline{p},F \overline{q′},F q′,F p); (F \overline{q′})^{∗}; (F \overline{p})_{∗}; (F p′)_{∗}; (F q)^{∗}; (F p)_{∗}; (F q′)^{∗}; (F p′)_{∗}; ,]  (13)

with Υ as defined in (10). Straightforward calculations show that these comparison cells satisfy the coherence conditions.

To check that these comparison cells satisfy the naturality conditions with respect to cells, consider the horizontally composable cells

\xymatrix{A_{1} \ar[d]_{a} & S_{1} \ar[l]_{q_{1}} \ar[d]_{s} \ar[r]^{p_{1}} & B_{1} \ar[d]_{b} & S′_{1} \ar[l]_{q′_{1}} \ar[d]_{s′} \ar[r]^{p′_{1}} & C_{1} \ar[d]^{c} \\ A_{2} & S_{2} \ar[l]^{q_{2}} \ar[r]_{p_{2}} & B_{2} & S′_{2} \ar[l]^{q′_{2}} \ar[r]_{p′_{2}} & C_{2}}

in 𝕊pan A with composition

\xymatrix{A_{1} \ar[d]_{a} & S_{1} ×_{B_{1}} S′_{1} \ar[l]_{q_{1}\overline{q′_{1}}} \ar[r]^{p′_{1}\overline{p_{1}}} \ar[d]_{b̃} & C_{1} \ar[d]^{c} \\ A_{2} & S_{2} ×_{B_{2}} S′_{2} \ar[l]^{q_{2}\overline{q′_{2}}} \ar[r]_{p′_{2}\overline{p_{2}}} & C_{2} .}

We need to show that

[figure: (F q_{1}\overline{q′_{1}})^{∗}; (F p′_{1}\overline{p_{1}})_{∗}; φ_{S_{1},S′_{1}}; (F q_{1}\overline{q′_{1}})^{∗}; F a; ζ; F s̃; ξ; (F p′_{1}\overline{p_{1}})_{∗}; F c; F a; (F q_{1})^{∗}; ζ; F s; (F p_{1})_{∗}; ξ; (F q′_{1})^{∗}; F b; ζ; F s′; (F p′_{1})_{∗}; ξ; F c; =; (F q_{2}\overline{q′_{2}})^{∗}; φ_{S_{2},S′_{2}}; (F p′_{2}\overline{p_{2}})_{∗}; (F q_{2})^{∗}; (F p_{2})_{∗}; (F q′_{2})^{∗}; (F p′_{2})_{∗}; (F q_{2})^{∗}; (F p_{2})_{∗}; (F q′_{2})^{∗}; (F p′_{2})_{∗}; .] (14)

By Lemma 3.14 we get that the left-hand side of this equation can be rewritten as

[figure: F (q_{1}\overline{q′_{1}})^{∗}; σ; F (p′_{1}\overline{p_{1}})_{∗}; τ; F (q_{1}\overline{q′_{1}})^{∗}; σ; F (p′_{1}\overline{p_{1}})_{∗}; τ; (F q_{1})^{∗}; Υ; (F \overline{q′_{1}})^{∗}; (F \overline{p_{1}})_{∗}; (F p′_{1})_{∗}; =; F a; (F q_{1})^{∗}; ζ; F s; (F \overline{q′_{1}})^{∗}; ζ; F s̃; (F \overline{p_{1}})_{∗}; ξ; F s′; (F p′_{1})_{∗}; ξ; F c; F a; (F q_{1})^{∗}; ζ; F s; (F p_{1})_{∗}; ξ; (F q′_{1})^{∗}; F b; ζ; F s′; (F p′_{1})_{∗}; ξ; F c; (F q_{2})^{∗}; Υ; (F \overline{q′_{2}})^{∗}; (F \overline{p_{2}})_{∗}; (F p′_{2})_{∗}; (F q_{2})^{∗}; (F p_{2})_{∗}; (F q′_{2})^{∗}; (F p′_{2})_{∗}; (F q_{2})^{∗}; (F p_{2})_{∗}; (F q′_{2})^{∗}; (F p′_{2})_{∗}; .]

It is now straightforward to check that

[figure: F (q_{1}\overline{q′_{1}})^{∗}; σ; F (q_{1}\overline{q′_{1}})^{∗}; F a; ζ; F s̃; F a; (F q_{1})^{∗}; ζ; F s; (F \overline{q′_{1}})^{∗}; ζ; F s̃; =; F (q_{2}\overline{q′_{2}})^{∗}; σ; (F q_{2})^{∗}; (F \overline{q′_{2}})^{∗}; (F q_{2})^{∗}; (F \overline{q′_{2}})^{∗}]

and

[figure: F (p′_{1}\overline{p_{1}})_{∗}; τ; F s̃; ξ; (F p′_{1}\overline{p_{1}})_{∗}; F c; F s̃; (F \overline{p_{1}})_{∗}; ξ; F s′; (F p′_{1})_{∗}; ξ; F c; =; τ; (F p′_{2}\overline{p_{2}})_{∗}; (F \overline{p_{2}})_{∗}; (F p′_{2})_{∗}; (F \overline{p_{2}})_{∗}; (F p′_{2})_{∗}]

and this gives us (14).

We conclude that F̃ is oplax normal. Also note that F̃ ∘ I_{A} = F, so composition with I_{A} is essentially surjective.

A natural transformation γ: F ⇒ F′: A ⇉ Vrt (𝔹) lifts to a vertical transformation γ̃: F̃ ⇒ F̃′: 𝕊pan A ⇉ 𝔹 defined as follows. For an object A in 𝕊pan A, γ̃_{A} is the vertical arrow γ_{A} in 𝔹. For a horizontal arrow \xymatrix{A & S \ar[l]_{q} \ar[r]^{p} & B}, the cell

\xymatrix{F̃A \ar[d]|{•}_{γ̃_{A}} \ar[r]_{γ̃_{S}}^{F̃(S)} & F̃B \ar[d]|{•}^{γ̃_{B}} \\ F̃′A \ar[r]_{F̃′(S)} & F̃′B}

is defined as

[figure: •; γ_{A}; (F q)^{∗}; ζ(F q,γ_{S},γ_{A},F′q); γ_{S}; •; (F p)_{∗}; ξ(F p,γ_{S},γ_{B},F′p); •; γ_{B}; (F′q)^{∗}; (F′p)_{∗}; .]

Note that if S is the identity span on A, then this diagram becomes the identity on γ_{A}. To show that γ̃ satisfies the (oplax) functoriality property with respect to horizontal

composition, let \xymatrix{A & S \ar[l]_{q} \ar[r]^{p} & B & S′ \ar[l]_{q′} \ar[r]^{p′} & C} be a composable pair of horizontal arrows

in 𝕊pan A with composition \xymatrix{A & T \ar[l]_{q\overline{q′}} \ar[r]^{p′\overline{p}} & C}. We need to show that

[figure: F (T); φ_{S,S′}; •; γ̃_{A}; γ̃_{T}; F (T); •; γ̃_{C}; F (S); •; γ̃_{A}; γ̃_{S}; •; γ̃_{B}; F (S′); γ̃_{S′}; •; γ̃_{C}; ≡; F′(T); φ′_{S,S′}; F′(S); F′(S′); F′(S); F′(S′)]  (15)

and the proof of this equation is completely analogous to that of (14).

The transformation γ̃ commutes trivially with the identity structure cells of F̃, since this functor preserves identities on the nose. Finally, note that γ̃ ∘ I_{A} = γ, so composition with I_{A} is full. To conclude that it is also faithful, note that γ̃ is completely determined by γ.

3.16. Remarks. Note that Theorem 3.15 states that 𝕊pan is a partial left biadjoint to the forgetful 2-functor

Vrt : Doub_{OplaxN} → Cat.

This is better than the bicategory version of the universal property of Span, Theorem 2.8, where we did not get a partial adjoint. That is because using the double category structure we were able to replace the condition on the morphism A → B, i.e., to be sinister, with a condition on the receiving double category 𝔹, i.e., to be gregarious. Or, put differently, when we pass from A to Span A we add right adjoints for the arrows of A, but this is not a completion. There are new arrows in Span A, which do not have right adjoints. However, passing from A to 𝕊pan A is a completion process. Every arrow in A gets a companion and conjoint and there are no new vertical arrows.

However, there are two points on which Theorem 3.15 is unsatisfactory. The first is the fact that as a left adjoint 𝕊pan is only defined for categories with pullbacks, not for all categories. The other issue is more subtle. Gregarious oplax morphisms of double categories do not preserve all horizontal composites. Generally, one would like morphisms to be structure preserving maps, so we would like to change the structure. Both of these issues will be resolved in Section 4.

3.17. Beck-Chevalley double categories. The most general equivalence of categories in the description of the universal property of the bicategory Span A in Theorem 2.8,

Bicat_{Greg}(Span A, B) ≃ Bicat_{Sinister}(A, B),  (16)

restricts to

Bicat_{Strong}(Span A, B) ≃ Bicat_{Beck}(A, B)

because strong morphisms from Span A to B correspond under (16) to sinister morphisms that satisfy the Beck-Chevalley condition.

So we may ask the question whether there is a similar restriction of the equivalence of categories in Theorem 3.15. In other words, what do strong morphisms 𝕊pan A → 𝔹 correspond to? It is clear that we will need to consider functors A → Vrt (𝔹) that preserve pullbacks, so we will need 𝔹 to have at least pullbacks of vertical arrows. This leads us to consider the following concept.

3.18. Definition. A double category

\xymatrix{A_{2} \ar[r]^{π_{1}} \ar[r]|{m} \ar[r]_{π_{2}} & A_{1} \ar[r]^{s} \ar[r]_{t} & A_{0} \ar[l]|{u}}

is said to have vertical pullbacks if this diagram lies within the category PBCat of categories with pullbacks and pullback-preserving functors between them.

A morphism F: 𝔸 → 𝔹 between categories with vertical pullbacks is said to preserve these pullbacks if its components F_{i}: A_{i} → B_{i} are pullback-preserving functors. When there is no confusion possible, we will also call such an F a pullback-preserving morphism.

Let us spell out what it means for a double category 𝔹 to have vertical pullbacks. If we represent 𝔹 by a diagram of categories and functors as in the definition above, the fact that u preserves pullbacks means that for every cospan of vertical arrows \xymatrix{A \ar[r]_{a} & C & B \ar[l]^{b}}

328  ROBERT DAWSON, ROBERT PARÉ, DORETTE PRONK

with pullback

\xymatrix{A ×_{C} B \ar[r]^{\overline{a}} \ar[d]_{\overline{b}} & B \ar[d]^{b} \\ A \ar[r]_{a} & C}

the corresponding square of unit cells

[figure: \overline{a}; \overline{b}; \overline{a}; \overline{b}; b; b; a; a]

is a pullback in B_{1}. As a consequence we obtain the following 2-dimensional universal property for pullback squares in B_{0}. For any horizontal arrow \xymatrix{X \ar[r]^{h} & Y} with cells

[figure: X; x_{A}; γ_{A}; h; Y; y_{A}; and; X; x_{B}; γ_{B}; h; Y; y_{B}; A; A; B; B]

such that

[figure: X; x_{A}; γ_{A}; h; Y; y_{A}; =; X; x_{B}; γ_{B}; h; Y; y_{B}; A; a; A; a; B; b; B; b; C; C; C; C]

there is a unique cell

[figure: X; h; x̃; γ̃; Y; ỹ; A ×_{C} B; A ×_{C} B]

such that 1\overline{b} · γ̃ = γ_{A} and 1\overline{a} · γ̃ = γ_{B}, as depicted in

[figure: Y; ỹ; y_{A}; y_{B}; X; h; x_{A}; x̃; γ̃; x_{B}; γ_{A}; A ×_{C} B; \overline{b}; \overline{a}; γ_{B}; B; b; A; a; C.]

Note that this is a generalization of the 2-dimensional universal property of pullbacks in 2-categories.

The fact that m preserves pullbacks means that one can horizontally compose two vertical pullback squares of cells and obtain again a vertical pullback square of cells. This is what was called “having a functorial choice of pullbacks” in [19].

3.19. Examples.

1. If A is a category with pullbacks, then the double category of pullback squares in A has vertical pullbacks.

2. If A is a category with pullbacks, then the double category 𝕊pan A has vertical pullbacks. The pullback of cells is depicted in the following diagram.

[figure: A_{1} ×_{A} A_{2}; _{\overline{a}_{2}}; _{\overline{a}_{1}}; S_{1} ×_{S} S_{2}; q̃; p̃; _{\overline{s}_{1}}; _{\overline{s}_{2}}; B_{1} ×_{B} B_{2}; _{\overline{b}_{2}}; _{\overline{b}_{1}}; A_{1}; a_{1}; S_{1}; q_{1}; p_{1}; s_{1}; B_{1}; b_{1}; A_{2}; a_{2}; S_{2}; s_{2}; q_{2}; p_{2}; B_{2}; b_{2}; A; T; q; p; B]  (17)

With a straightforward diagram chase using the universal properties of all pullbacks involved the reader may verify that all the structure maps are pullback-preserving.

3. Let V be a monoidal category with sums (that distribute over tensor) and pullbacks. Then there is a double category V -𝕄at with sets as objects, functions as vertical arrows, and V -matrices [V_{ij}]: I → J as horizontal arrows. The cells are of the form

[figure: I; f; [V_{ii′}]; [α_{ii′}]; I′; f′; K; [W_{kk′}]; K′]

where the α_{ii′}: V_{ii′} → W_{f(i)f′(i′)} are arrows in V. To define the vertical pullback of squares

\xymatrix{I \ar[dd]_{f} \ar[r]_{[α_{ii′}]}^{[V_{ii′}]} & I′ \ar[dd]^{f′} &  & J \ar[dd]_{g} \ar[r]_{[β_{jj′}]}^{[W_{jj′}]} & J′ \ar[dd]^{g′} \\  &  & and \\ K \ar[r]_{[X_{kk′}]} & K′ &  & K \ar[r]_{[X_{kk′}]} & K′}

let P = I ×_{K} J and P′ = I′ ×_{K′} J′ in Sets. Moreover, for (i, j) ∈ P and (i′, j′) ∈ P′, let

\xymatrix{Y_{(i,j),(i′,j′)} \ar[r]^{^{\overline{α}^{(i,j),(i′,j′)}}} \ar[d]_{_{\overline{β}_{(i,j),(i′,j′)}}} & W_{jj′} \ar[d]^{β_{jj′}} \\ V_{ii′} \ar[r]_{α_{ii′}} & X_{kk′}}

(where k = f(i) and k = g(j), and k′ = f′(i′) and k′ = g′(j′)) be a pullback in V. Then the pullback of the cells [α_{ii′}] and [β_{jj′}] is

[figure: P; π_{2}; π_{1}; [Y_{(i,j),(i′,j′)}]; [\overline{β}_{(i,j),(i′,j′)}]; [\overline{α}_{(i,j),(i′,j′)}]; P′; π_{2}; π_{1}; I; [V_{ii′}]; f; [α_{ii′}]; I′; f′; J; g; [W_{jj′}]; J′; g′; [β_{jj′}]; K; [X_{kk′}]; K′ .]

We see now that the horizontal composition of vertical pullback squares of cells will only give a pullback square of cells if the sums and tensors distribute over the pullbacks.

Moreover, this double category is gregarious. For a vertical arrow f: I → K, we define the companion horizontal arrow as f_{∗}: I → K with (f_{∗})_{ik} = I if f(i) = k and (f_{∗})_{ik} = 0 otherwise. The binding cells are defined in the obvious way. For example, χ_{ik} = 1_{I} if f(i) = k and χ_{ik} = 1_{0} otherwise. The conjoint f^{∗}: K → I is defined by (f^{∗})_{ki} = I if f(i) = k and (f^{∗})_{ki} = 0 otherwise.

The fact that the vertical pullback of cells in 𝕊pan A is formed as in (17) generalizes to the following result for gregarious double categories.

3.20. Proposition. Let 𝔸 be a gregarious double category. Then the functor Ξ: A2_{0} → A_{1}, defined in (5) and (6), and the functor Z: A2_{0} → A_{1}, defined in (7) and (8), preserve all limits that the functor U: A_{0} → A_{1} (which sends objects and vertical arrows to their horizontal identities) preserves.

Proof. We will show this for Ξ. The proof for Z is its dual. Let P(𝔸) be the category whose objects are diagrams of the form

\xymatrix{A \ar[r]|{•}^{x} & B \ar[r]^{f} & C}

in 𝔸, and arrows are commutative diagrams of the form

[figure: A; x; •; •; v_{1}; B; f; •; v_{2}; α; C; •; v_{3}; A′; x′; •; B′; f′; C′]

in 𝔸.

The functor Ξ factors through P(𝔸) in the following way:

\xymatrix{A2_{0} \ar[rr]^{Ξ} \ar[dr]_{L} &  & A_{1} \\  & P(𝔸) \ar[ur]_{Γ}}

where L is defined by

[figure: (; A; •; x; B; ); ↦; (; A; •; x; B; B; )]

[figure: (; A; •; x; •; u; B; •; v; A′; •; x′; B′; ); ↦; (; A; •; x; •; u; B; •; v; 1_{v}; B; •; v; A′; •; x′; B′; B′; )]

and Γ is defined by

[figure: (; A; •; x; B; f; C; ); ↦; (; A; •; fx_{∗}; C; )]

( [figure: A; x; •; •; v_{1}; B; f; •; v_{2}; α; C; •; v_{3}; A′; x′; •; B′; f′; C′]) ↦ ( [figure: A; fx_{∗}; •; v_{1}; αξ(x,v_{1},v_{2},x′); C; •; v_{3}; A′; f′(x′)_{∗}; C′]) .

A straightforward calculation shows that L preserves all limits that U preserves.

We will now show that Γ has a left adjoint and therefore preserves all limits. The left adjoint is the functor Δ: A_{1} → P(𝔸) defined by

[figure: (; A; f; B; ); ↦; (; A; A; f; B; )]

( \xymatrix{A \ar[d]|{•}_{u} \ar[r]_{α}^{f} & B \ar[d]|{•}^{v} \\ A′ \ar[r]_{f′} & B′}) ↦ ( [figure: A; •; u; A; •; u; α; f; B; •; v; A′; A′; f′; B′]) .

Note that the composite ΔΓ: P(𝔸) → P(𝔸) is given by

[figure: (; A; •; x; B; f; C; ); ↦; (; A; A; fx_{∗}; B; )]

( \xymatrix{A \ar[d]|{•}_{v_{1}} \ar[r]|{•}^{x} & B \ar[d]|{•}_{v_{2}} \ar[r]_{α}^{f} & C \ar[d]|{•}^{v_{3}} \\ A′ \ar[r]|{•}^{x′} & B′ \ar[r]_{f′} & C′} ) ↦ ( [figure: A; •; v_{1}; A; •; v_{1}; αξ(x,v_{1},v_{2},x′); fv_{∗}; C; •; v_{3}; A′; A′; f′(x′)_{∗}; B′] ) .

We define the counit to be the natural transformation ε: ΔΓ ⇒ 1_{P(𝔸)} by

ε_{(x,f)} = [figure: •; x; fx_{∗}; id_{f} χ_{x}; •; x; f]

To check that this is indeed natural, let

[figure: A; x; •; •; v_{1}; B; f; •; v_{2}; α; C; •; v_{3}; A′; x′; •; B′; f′; C′]

be any arrow in P(𝔸). Then we need to show that

[figure: •; x; fx_{∗}; id_{f} χ_{x}; •; v_{1}; •; v_{1}; fx_{∗}; αξ(x,v_{1},v_{2},x′); •; v_{3}; •; v_{1}; •; x; •; v_{2}; f; α; •; v_{3}; =; •; x′; f′(x′)_{∗}; id_{f′}χ_{x′}; •; x′; f′; •; x′; f′; ,]

i.e., that

[figure: •; v_{1}; •; v_{2}x; fx_{∗}; αχ_{x}; •; v_{3}; =; •; v_{1}; •; v_{1}; fx_{∗}; (id_{f′}χ_{x′})·(αξ(x,v_{1},v_{2},x′)); •; v_{3}; •; x′; f′; •; x′; f′; ,]

and this follows immediately from the definition of ξ(x, v_{1}, v_{2}, x′) and one of the binding cell identities for x′.

Note that we can take (id_{A})_{∗} = 1_{A} and ξ(id_{A}, u, u, id_{B}) = 1_{u}, so it follows that ΓΔ = 1_{A_{1}}, and we may take the unit of the adjunction to be the identity transformation. This completes the proof of this proposition.

3.21. Corollary. Let 𝔹 be a gregarious double category with vertical pullbacks. Any commutative cube of vertical arrows

[figure: A_{1} ×_{A} A_{2}; \overline{a_{1}}; ṽ; \overline{a_{2}}; B_{1} ×_{B} B_{2}; \overline{b_{1}}; \overline{b_{2}}; A_{1}; a_{1}; v_{1}; B_{1}; b_{1}; A_{2}; a_{2}; v_{2}; B_{2}; b_{2}; A; v; B]

where the left and right faces are pullback squares, gives rise to the following vertical pullbacks of cells:

[figure: A_{1} ×_{A} A_{2}; ξ(ṽ,\overline{a_{2}},\overline{b_{2}},v_{1}); \overline{a_{1}}; (ṽ)_{∗}; \overline{a_{2}}; B_{1} ×_{B} B_{2}; \overline{b_{1}}; \overline{b_{2}}; A_{1}; a_{1}; (v_{1})_{∗}; ξ(v_{1},a_{1},b_{1},v); B_{1}; b_{1}; A_{2}; a_{2}; (v_{2})_{∗}; ξ(ṽ,\overline{a_{1}},\overline{b_{1}},v_{2}); B_{2}; b_{2}; A; v_{∗}; ξ(v_{2},a_{2},b_{2},v); B]  (18)

and

[figure: B_{1} ×_{B} B_{2}; ζ(ṽ,\overline{b_{2}},\overline{a_{2}},v_{1}); \overline{b_{1}}; \overline{b_{2}}; A_{1} ×_{A} A_{2}; (ṽ)^{∗}; \overline{a_{1}}; \overline{a_{2}}; B_{1}; ζ(v_{1},b_{1},a_{1},v); b_{1}; A_{1}; a_{1}; (v_{1})^{∗}; B_{2}; b_{2}; ζ(ṽ,\overline{b_{1}},\overline{a_{1}},v_{2}); A_{2}; a_{2}; (v_{2})^{∗}; B; ζ(v_{2},b_{2},a_{2},v); A .; v^{∗}]  (19)

We would like to restrict the equivalence in Theorem 3.15 in such a way that we obtain pullback-preserving strong morphisms 𝕊pan A → 𝔹 as objects on the left-hand side of the equivalence.

Note that we cannot put the Beck-Chevalley condition on the functors from A into Vrt (𝔹), since Vrt (𝔹) doesn’t contain the information about the companions and conjoints in 𝔹. However, just as with the notion of gregariousness, we can make the Beck-Chevalley condition a property of gregarious double categories rather than a property of the morphisms between them.

3.22. Definition. A Beck-Chevalley double category is a gregarious double category which has vertical pullbacks and for which the following version of the Beck-Chevalley condition holds: For each pullback square of vertical arrows

[figure: k; h; g; f]

the induced cell

[figure: h^{∗}; β_{h}; h; k; χ_{k}; k_{∗}; Υ(k, h, g, f); =; ψ_{f}; f; g; α_{g}; f_{∗}; g^{∗}]

(as defined in (10)) is vertically invertible. We will call the cells of the form Υ(k, h, g, f) Beck-Chevalley cells.

We write BCDoub for the 2-category of Beck-Chevalley double categories with strong pullback-preserving morphisms and vertical transformations.

3.23. Examples.

1. For any category A with pullbacks, the double category 𝕊pan(A) is Beck-Chevalley.

2. The double category ℝel with sets as objects, functions as vertical arrows and relations as horizontal arrows is a Beck-Chevalley double category.

3. In V -𝕄at (with V as in Example 3.19 (3)) let

\xymatrix{P \ar[r]|{•}^{q} \ar[d]|{•}_{p} & J \ar[d]|{•}^{g} \\ I \ar[r]|{•}_{f} & K}

be a vertical pullback square. Then

(g^{∗}f_{∗})_{ij} = ∑_{k∈K}(f_{∗})_{ik} ⊗ (g^{∗})_{kj} = \{ \begin{matrix}I & if f(i) = g(j) \\ 0 & otherwise \end{matrix}

and

(q_{∗}p^{∗})_{ij} = ∑_{\begin{matrix}r,s \\ f(r)=g(s) \end{matrix}}(q_{∗})_{i,(r,s)} ⊗ (p^{∗})_{(r,s),j} = \{ \begin{matrix}I & if f(i) = g(j) \\ 0 & otherwise \end{matrix}

and Υ(q, p, g, f) is the vertical identity cell. So V -𝕄at is a Beck-Chevalley double category (if V satisfies the conditions in Example 3.19 (3)).

We are now ready to state the restriction of the equivalence in Theorem 3.15 to pullback-preserving strong morphisms of double categories.

3.24. Theorem. Let A be a category with pullbacks and 𝔹 be a Beck-Chevalley double category. Then composing with the inclusion A ↪ 𝕊pan A gives an equivalence of categories between the category of strong pullback-preserving morphisms 𝕊pan A → 𝔹 with vertical transformations, and the category of pullback-preserving functors A → Vrt (𝔹) with natural transformations,

BCDoub (𝕊pan A, 𝔹) ≃ PBCat (A, Vrt (𝔹)).

Proof. It is clear that for any pullback-preserving strong morphism F: 𝕊pan A → 𝔹, the composition with the inclusion A ↪ 𝕊pan A induces a pullback-preserving functor of categories A → Vrt 𝔹.

To show that this functor BCDoub(𝕊pan A, 𝔹) → PBCat (A, Vrt (𝔹)) is essentially surjective, let F: A → Vrt 𝔹 be a pullback-preserving functor and let F̃: 𝕊pan A → 𝔹 be the lifting constructed in the proof of Theorem 3.15.

Since 𝔹 is a Beck-Chevalley double category and F preserves pullbacks, the cell Υ(F \overline{p}, F \overline{q′}, F q′, F p) is vertically invertible for every pullback diagram

\xymatrix{S ×_{B} S′ \ar[d]_{\overline{q′}} \ar[r]^{\overline{p}} & S′ \ar[d]^{q′} \\ S \ar[r]_{p} & B}

in A. Diagram (13) shows that this implies that the comparison cells for F̃ are all isomorphisms (since σ and τ are vertical isomorphisms), so F̃ is a strong morphism.

We also need to show that F̃ preserves all vertical pullbacks. It preserves pullbacks of vertical arrows by the assumption that F preserves pullbacks. We will consider the image of the pullback diagram (17) under F̃. To make our notation more manageable, let ζ_{i} = ζ(F q_{i}, F s_{i}, F a_{i}, F q), and ξ_{i} = ξ(F p_{i}, F s_{i}, F b_{i}, F p) for i = 1, 2. Also, let \overline{ζ}_{1} = ζ(F q̃, F \overline{s}_{1}, F \overline{a}_{1}, F q_{2}), \overline{ζ}_{2} = ζ(F q̃, F \overline{s}_{2}, F \overline{a}_{2}, F q_{1}), \overline{ξ}_{1} = ξ(F p̃, F \overline{s}_{1}, F \overline{b}_{1}, F p_{2}), and \overline{ξ}_{2} = ξ(F p̃, F \overline{s}_{2}, F \overline{b}_{2}, F p_{1}). Now we need to show that the vertical square of cells

[figure: ^{\overline{ξ}^{2}\overline{ζ}^{2}}; _{\overline{ξ}_{1}\overline{ζ}_{1}}; ξ_{1}ζ_{1}; ξ_{2}ζ_{2}]  (20)

is a pullback square. Corollary 3.21 gives us that the squares

[figure: ^{\overline{ζ}^{2}}; _{\overline{ζ}_{1}}; ζ_{1}; and; ^{\overline{ξ}^{2}}; _{\overline{ξ}_{1}}; ξ_{1}; ζ_{2}; ξ_{2}]

are vertical pullbacks. The fact that (20) is a pullback follows now from the fact that the horizontal composition of two vertical pullback squares of cells is again a vertical pullback square.

3.25. Functoriality. We saw in Section 2.9 that the bicategory version of Span is not a 2-functor. The double category version, however, is. As implied by Theorem 3.15 and Theorem 3.24, the 𝕊pan construction should give bifunctors from pbCat to Doub_{OplaxN} and from PBCat to BCDoub. In fact, more is true.

3.26. Proposition. The Span construction gives a 2-functor 𝕊pan: pbCat → Doub_{OplaxN} which restricts to a 2-functor 𝕊pan: PBCat → BCDoub.

One consequence of 2-functoriality is that 𝕊pan preserves adjunctions. This will give us completeness results for double categories of the form 𝕊pan A. There are various equivalent ways of discussing limits and colimits in double categories (see for example [19]). Definition 3.28 below is best suited to our purpose.

3.27. Proposition. The 2-categories pbCat, PBCat, Doub_{OplaxN}, and BCDoub are cotensored and the 2-functors

\xymatrix{PBCat \ar[d] \ar[r]^{𝕊pan} & BCDoub \ar[d] \\ pbCat \ar[r]^{𝕊pan} & _{Doub_{OplaxN}}}

preserve cotensors.

Proof. If A is a category with pullbacks and I an arbitrary category, the functor category A^{I} also has pullbacks, and cartesian closedness of Cat gives the cotensor property for pbCat:

I ⟶ pbCat (B, A) ; B ⟶ A^{I} in pbCat

The bijection restricts to a bijection

I ⟶ PBCat (B, A) ; B ⟶ A^{I} in PBCat

This is quite general and would work for any class of limits or colimits in place of pullbacks. Given a weak double category 𝔸, it can be considered a weak category object in Cat,

\xymatrix{A_{2} \ar[r] \ar[r] \ar[r] & A_{1} \ar[r] \ar[r] & A_{0}. \ar[l]}

Since ( )^{I} is a 2-functor, it follows that

\xymatrix{A_{2}^{I} \ar[r] \ar[r] \ar[r] & A_{1}^{I} \ar[r] \ar[r] & A_{0}^{I}. \ar[l]}

is again a weak double category object, which we shall denote by 𝔸^{I}. A basic calculus of adjoints shows that 𝔸^{I} is indeed the cotensor of 𝔸 by I in Doub_{OplaxN}, i.e., we have an isomorphism of categories

Doub_{OplaxN}(𝔹, 𝔸^{I}) ≅ Cat(I, Doub_{OplaxN}(𝔹, 𝔸)).

This isomorphism is easily seen to restrict to an isomorphism

BCDoub(𝔹, 𝔸^{I}) ≅ Cat(I, BCDoub(𝔹, 𝔸)).

It is clear from the constructions that cotensors are preserved by the 2-functors 𝕊pan and ‘inclusion’.

3.28. Definition. A double category 𝔸 has oplax I-colimits if the diagonal Δ: 𝔸 → 𝔸^{I} has a left adjoint in the 2-category Doub_{OplaxN}. The I-colimits are strong if Δ has a left adjoint in Doub. A dual definition gives lax and strong I-limits.

3.29. Theorem. If A has I-colimits then 𝕊pan A has oplax I-colimits. If the I-colimits commute with pullbacks in A then 𝕊pan A has strong I-colimits. If A has I-limits, then 𝕊pan A has strong I-limits.

Proof. If A has I-colimits we obtain a functor lim_{→}: A^{I} → A which is left adjoint to

Δ: A → A^{I} in pbCat and 𝕊pan: pbCat → Doub_{OplaxN} preserves this adjunction. As 𝕊pan(A^{I}) ≅ 𝕊pan(A)^{I} we see that 𝕊pan A has oplax I-colimits.

If I-colimits commute with pullbacks in A, the adjunction lim_{→} ⊣ Δ lives in PBCat and 𝕊pan maps PBCat to BCDoub. The arrows of BCDoub are strong morphisms, so 𝕊pan A has strong I-colimits.

The result for I-limits follows from the fact that the adjunction Δ ⊣ lim_{←} is always in PBCat.

3.30. Remarks.

1. This proposition extends our result about pullbacks in the previous subsection to arbitrary small limits. Another important special case occurs when A has products. It is well-known that products in A give rise to a tensor structure on Span(A). However, in 𝕊pan(A) they give rise to vertical products with a universal property.

2. The fact that 𝕊pan of a left adjoint in Cat only gives rise to an oplax morphism of double categories fits in the general theory of adjunctions between double categories as developed by Grandis and Paré in [18]; generally, left adjoints are oplax whereas right adjoints are lax.

3. There are important cases when the colimit functor preserves pullbacks and thus these vertical limits in 𝕊pan A will be strong. A category A is lextensive if it has finite limits and finite coproducts with the property that the coproduct functor

+: A/I × A/J → A/(I + J)

is an equivalence (and also A/0 ≃ 1). Thus, + preserves pullbacks so that 𝕊pan A has strong finite coproducts.

In a similar vein, if A is a Grothendieck topos, the double category 𝕊pan A will have arbitrary strong coproducts. In fact, disjoint and universal coproducts suffice.

If A is locally finitely presentable, then filtered colimits commute with finite limits, so 𝕊pan A has strong filtered colimits.

3.31. Equipments. A different way to make the Span construction 2-functorial was presented in [7]. This paper makes a very careful study of the types of morphisms one may want to consider between categories of spans. The spn construction of [7] produces what is called a pointed, starred equipment. An equipment has objects, scalar arrows, vector arrows, and vector transformations. The objects and vector arrows form a category, and for any pair of objects A and B, there is a category with the vector arrows from A to B as objects and the vector transformations between them as morphisms. Finally, the scalar arrows act on the vector arrows and vector transformations from both the left and the right in such a way that if k: K → K′ and l: L → L′ are scalar arrows, μ, ν: \xymatrix{K \ar[r]|{+} & L} are vector arrows, and and Φ: μ → ν is a vector transformation, then there are vector transformations Φk: μk → νk and lΦ: lμ → lν, and these actions are functorial in Φ, strictly unitary, and coherently associative in all three possible senses.

An equipment is called starred if for each scalar k the action (−)k has a left adjoint, denoted by (−)k^{∗} and for each scalar l, the action l(−) has a right adjoint, denoted by l^{∗}(−), and these actions need to satisfy a Beck-Chevalley type of condition.

An equipment is called pointed if it comes a with a family of distinguished vector arrows [figure: ι_{K}: K; +; K] for each object K and an invertible vector transformation \xymatrix{fι_{K} \ar[r]^{≃} & ι_{L}f}

for each scalar f: K → L, such that ( \xymatrix{1ι_{K} \ar[r]^{≃} & ι_{K}1} ) = 1_{ι_{K}}.

In order to compare gregarious double categories to equipments, we note that analogously to the case of adjunctions, companions can also be characterized by the following universal properties:

3.32. Lemma. For a vertical arrow \xymatrix{A \ar[r]|{•}^{v} & B} in a double category the following are equivalent:

1. v is a companion;

2. For each horizontal arrow \xymatrix{X \ar[r]^{f} & A}, there is a horizontal arrow \xymatrix{X \ar[r]^{v∗f} & B} and a

cell

[figure: X; f; ψ_{v,f}; A; •; v; X; v∗f; B]

which is initial in the sense that for each cell α as in (21) there is a unique cell \overline{α}:

[figure: •; x; α; f; •; v; ψ_{v,f}; f; •; v; •; w; =; •; x; v∗f; \overline{α}; •; w; y; y]  (21)

3. For each horizontal arrow \xymatrix{B \ar[r]^{g} & Y} there is a horizontal arrow \xymatrix{A \ar[r]^{g∗v} & Y} and a

cell

[figure: A; •; v; g∗v; χ_{v,g}; Y; B; g; Y]

which is terminal in the the sense that for each cell β as in the following diagram there is a unique cell β̃:

[figure: •; u; z; β; •; x; •; u; z; β̃; •; x; •; v; =; •; v; g∗v; χ_{v,g}; g; g]

Proof. To show that 1 implies 2, let v have a horizontal companion v_{∗} with binding cells ψ_{v} and χ_{v}, and let f be as in 2. Then take v ∗ f := v_{∗}f and the cell

[figure: f; ψ_{v,f} :=; f; ψ_{v}; •; v; f; v_{∗}; .]

For any cell α as in (21) the induced cell \overline{α} is

[figure: •; x; α; f; v_{∗}; χ_{v}; •; v; •; w; •; w; y; .]

It is clear that this is the unique cell that fits in the diagram (21).

To show that 2 implies 1, let v be a vertical arrow with the property described in 2. Define the horizontal companion of v to be v_{∗} := v ∗ 1_{A}, with binding cells ψ_{v} = v and χ_{v} = \overline{1_{v}}, where \overline{1_{v}} is the unique cell such that

[figure: •; v; 1_{v}; •; v; ψ_{v,1_{A}}; •; v; =; •; v; v∗1_{A}; \overline{1_{v}}; ,]

as in (21). This establishes right away that these binding cells satisfy the second binding cell identity in Definition 3.4. To prove that they also satisfy the first one, note that

[figure: ψ_{v,1A}; •; v; ψ_{v,1A}; •; v; ψ_{v,1A}; •; v; ψ_{v,1A}; •; v; v∗1_{A}; \overline{1_{v}}; =; =; v∗1_{A}; _{id_{v∗1A}}; v∗1_{A}; v∗1_{A}; v∗1_{A}; .]

By the uniqueness of \overline{α} in (21) we conclude that \overline{1_{v}}ψ_{v,1_{A}} = id_{v∗1_{A}}.

The proof that 1 is equivalent to 3 is dual to the one that 1 is equivalent to 2 and left for the reader.

By duality, conjoints can be characterized in the following way:

3.33. Lemma. For a vertical arrow \xymatrix{A \ar[r]|{•}^{v} & B} in a double category the following are equivalent:

1. v is a conjoint;

2. For each horizontal arrow \xymatrix{A \ar[r]^{f} & X}, there is a horizontal arrow \xymatrix{B \ar[r]^{f⋆v} & X} and a

cell

[figure: A; f; •; v; α_{v,f}; X; B; f⋆v; X]

which is initial in the sense that for each cell ζ as in Figure 1 there is a unique cell \overline{ζ} as in Figure 1.

[figure: •; v; ζ; f; •; x; •; v; α_{v,f}; f; •; w; =; •; w; f⋆v; \overline{ζ}; •; x; y; y]

Figure 1: For any cell ζ there is an induced cell \overline{ζ}.

3. For each horizontal arrow \xymatrix{Y \ar[r]^{g} & B} there is a horizontal arrow \xymatrix{Y \ar[r]^{v⋆g} & A} and a

cell

[figure: Y; v⋆g; β_{v,g}; A; •; v; Y; g; B]

which is terminal in the the sense that for each cell θ as in Figure 2 there is a unique cell θ̃ as in Figure 2.

[figure: •; x; z; θ; •; u; •; x; z; θ̃; •; u; •; v; =; v⋆g; β_{v,g}; •; v; g; g]

Figure 2: For any cell θ there is an induced cell θ̃.

Given a double category with companions, one can construct a pointed equipment by using the same objects, defining the scalar arrows to be the vertical arrows of the double category, the vector arrows to be the horizontal arrows of the double category, and the vector transformations correspond to special cells (where the horizontal domain and codomain are identity arrows). Composition is defined as in the double category, and the action of the scalar arrows is defined using companions and horizontal composition. With the notation as above, lμ := l_{∗}μ (where the concatenation on the left-hand side stands for horizontal composition), μk := μk_{∗}, and lΦ := id_{l_{∗}}Φ and Φk := Φid_{k_{∗}}. The pointing is defined by the horizontal identity arrows (the universal properties in Lemma 3.32 imply the existence of a canonical vertically invertible special cell from v ∗ 1_{A} to 1_{B} ∗ v for any vertical arrow \xymatrix{v: A \ar[r]|{•} & B} ).

For a Beck-Chevalley double category this construction can be extended to obtain a starred pointed equipment. Oplax normal morphisms between double categories give rise to right homomorphisms of equipments under this construction.

However, since equipments do not have a notion of horizontal composition, we cannot reverse this operation to create a double category out of an equipment. The best we could hope for is an oplax double category and we will return to this question in the next section.

4. Spans without pullbacks

Notice that Theorem 2.8 gives a characterization of Span A that does not refer to pullbacks, and Theorem 3.15 gives a similar characterization for 𝕊pan A. So we might use these universal properties to define Span A and 𝕊pan A for general categories. This would be useful as we could then apply the results of [12] to Span A and 𝕊pan A for further study of the Π_{2}-construction.

Of course, pullbacks are necessary in the composition of spans and so we take the drastic step of discarding composition. There are several justifications for such an action. The most immediate is that without pullbacks the composites just are not there, but there is sufficient extra structure at the 2-dimensional level to salvage a good part of what composition does for us. Another consideration is that the universal property expressed in Theorem 2.8 refers to oplax morphisms out of Span A, and if the relevant morphisms do not preserve composition, we should not require it to be there. Of course, oplax morphisms do involve composition and the 2-dimensional structure referred to before is exactly what is needed to express this.

Along the same lines, for categories A and B with pullbacks and a functor F: A → B that does not preserve them, Span(F): Span A → Span B is only oplax. Clearly, we are in need of oplax bicategories or some such concept. There exist two notions which might serve in this capacity, the dual of Hermida’s bicategory with several objects [21] and the dual of Leinster’s fc-multicategory [26]. Both Hermida and Leinster use spans in a category without pullbacks as a basic example. It is our contention that the second notion is the most useful one, at least for the work that follows. Leinster’s fc-multicategories can also be viewed as Burroni’s 𝕋-multicategories in Set^{•⇉•}, where 𝕋 is the free category monad [5], or as the 2-dimensional truncation of Hermida’s multicategories in ω-graphs, introduced in Section 10.2 of [22]. Hermida defines a notion of representability for composites which differs a bit from ours, as we will discuss in Section 4.6.

In [12], we called fc-multicategories lax double categories. We showed that lax morphisms of double categories are the Kleisli morphisms for the comonad ℙath on the 2category Doub of double categories with double functors and vertical transformations, and that the Eilenberg-Moore algebras are precisely the lax double categories. (The anonymous referee has made the interesting suggestion that this result follows from Theorem 6.1 in [22].) We took this fact as justification that lax double categories are the right structure for discussing lax morphisms.

4.1. Oplax double categories. Recall from [12] that an oplax double category has objects, horizontal and vertical arrows, and cells. There are two differences with double categories. The first is that there is no composition given for horizontal arrows. This is somewhat compensated for by the fact that cells are more general. They have vertical arrows as horizontal domain and codomain and although the vertical domain is a horizontal arrow, the vertical codomain is a finite string of compatible horizontal arrows (possibly of length 0). So a cell is of the form

[figure: A; •; v; α; f; B; •; w; C_{0}; g_{1}; C_{1}; g_{2}; C_{2}; g_{3}; g_{n}; C_{n}.]  (22)

Vertical arrows compose and form a category. There are identity cells

[figure: A; _{id_{f}}; f; B; A; f; B]

and cells compose vertically in the way that functions of several variables do. To be precise, given cells

[figure: C_{i−1}; g_{i}; •; x_{i−1}; β_{i}; C_{i}; •; x_{i}; D_{i0}; D_{i1}; D_{im_{i}}]

for i = 1, …, n, we can compose them with α in (22) to obtain a cell Note that there are n + 1 vertical arrows x_{i} and n cells β_{i} involved in such a composite, so when n = 0 our composite looks like

[figure: A; f; (β_{1},...,β_{n})·α; x_{0}·v; •; B; •; x_{n}·w; D_{10}; D_{11}; D_{nm_{n}}.]

[figure: A; •; v; f; α; B; •; w; C_{0}; •; x_{0}; D_{10} .]

This composition is unitary and associative in the obvious sense.

A morphism of oplax double categories is a map that preserves all the structure (on the nose).

Every double category 𝔸 gives rise to a canonical oplax double category 𝕆plax 𝔸 with the same objects and horizontal and vertical arrows, but where a cell α as in (22) corresponds to a cell

[figure: A; •; v; f; α; B; •; w; C_{0}; g_{n}g_{n−1}···g_{1}; C_{n}]

in 𝔸. Then morphisms of oplax double categories 𝕆plax 𝔸 → 𝕆plax 𝔹 correspond precisely to oplax morphisms 𝔸 → 𝔹. All of this is explained in full detail in [12].

An oplax bicategory is an oplax double category where all the vertical arrows are identities. Lax double categories and lax bicategories are defined dually.

4.2. Definition. Let A be an arbitrary category. Then 𝕊pan A is the oplax double category whose objects and vertical arrows are the objects and arrows of A, respectively. The horizontal arrows of 𝕊pan A are spans in A and its cells are commutative diagrams of the form

[figure: A; f; S; p; p′; x_{1}; x_{2}; x_{n}; A′; f′; B_{0}; T_{1}; B_{1}; T_{2}; B_{2}; B_{n−1}; T_{n}; B_{n}]  (23)

4.3. Proposition. With the above arrows and cells, 𝕊pan A can be made into an oplax double category.

Proof. Identities are given by

[figure: A; S; A′; A; S; A′]

and composition is defined as follows. Suppose we are given n cells

[figure: B_{i−1}; g_{i−1}; T_{i}; y_{i1}; y_{i2}; y_{im_{i}}; B_{i}; g_{i}; C_{i,0}; U_{i1}; C_{i,1}; U_{i2}; U_{im_{i}}; C_{i,m_{i}},]

with C_{i,0} = C_{i−1,m_{i−1}}, the composite with the cell (23) above is

[figure: A; g_{0}f; S; y_{ij}x_{i}; A′; g_{n}f′; C_{10}; U_{ij}; C_{n,m_{n}}.]

Verification of associativity is straightforward.

Let F: A → B be a functor. We get a morphism of oplax double categories,

𝕊pan(F): 𝕊pan A → 𝕊pan B,

simply by applying F to each of the components of arrows or cells. So, for example, 𝕊pan(F) applied to the span A ⟵^{p} S ⟶^{p′} A′ gives the span F A ⟵^{F p} F S ⟶^{F p′} F A′, and so on.

For a natural transformation t: F → G we get a vertical transformation

𝕊pan t: 𝕊pan F → 𝕊pan G,

defined by

\xymatrix{ & F A \ar[dd]|{•}^{t_{A}} \\ (𝕊pan t)_{A} := \\  & GA \\  & F A \ar[dd]|{•}_{t_{A}} & F S \ar[l]_{F p} \ar[dd]|{•}_{t_{S}} \ar[r]^{F p′} & F A′ \ar[dd]|{•}^{t_{A′}} \\ (𝕊pan t)_{S} := \\  & GA & GS \ar[l]^{Gp} \ar[r]_{Gp′} & GA.}

We have the following result.

4.4. Proposition. Let Oplax be the 2-category of oplax double categories, oplax morphisms and vertical transformations. Then 𝕊pan: Cat → Oplax is a locally fully faithful 2-functor.

Proof. The fact that 𝕊pan is a 2-functor can be verified by a straightforward calculation. (See [12], following Theorem 2.1 for the definition of vertical transformation.) Local faithfulness is also obvious. To check that 𝕊pan is locally full, let τ: 𝕊pan F → 𝕊pan G be any vertical transformation, where F, G: A ⇉ B are functors. Let τ_{(p,q)} be of the form

\xymatrix{F A \ar[d]_{τ_{A}} & F S \ar[l]_{F p} \ar[r]^{F q} \ar[d]_{τ_{(p,q)}} & F B \ar[d]^{τ_{B}} \\ GA & GS \ar[l]^{Gp} \ar[r]_{Gq} & GB}

We want to show that τ_{(p,q)} = τ_{S}. First note that τ_{(1_{S},1_{S})} = τ_{S}. Now consider the 2-cell

\xymatrix{S \ar[d]_{p} & S \ar[l]_{1} \ar[r]^{1} \ar[d]_{1} & S \ar[d]^{q} \\ A & S \ar[l]^{p} \ar[r]_{q} & B}

in 𝕊pan A. By vertical naturality of τ, we have that

\xymatrix{F S \ar[d]_{F p} & F S \ar[l]_{F 1} \ar[d]_{F 1} \ar[r]^{F 1} & F S \ar[d]^{F q} &  & F S \ar[d]_{τ_{S}} & F S \ar[l]_{F 1} \ar[r]^{F 1} \ar[d]_{τ_{(1,1)}=τ_{S}} & F S \ar[d]^{τ_{S}} \\ F \ar[d]_{τ_{A}} & F S \ar[d]_{τ_{(p,q)}} \ar[r]^{F q} & F B \ar[d]^{τ_{B}} & = & GS \ar[d]_{Gp} & GS \ar[l]^{G1} \ar[r]_{G1} \ar[d]_{G1} & GS \ar[d]^{Gq} \\ GA & GS \ar[l]^{Gp} \ar[r]_{Gq} & GB &  & GA & GS \ar[l]^{Gp} \ar[r]_{Gq} & GB .}

We conclude that τ_{(p,q)} = τ_{S}, as required.

4.5. Proposition. The inclusion A ↪ 𝕊pan A as vertical arrows is locally full and faithful.

Proof. Given f, g: A ⇉ B in A, there is at most one cell

[figure: A; f; A; A; g; B; B; B]

and there is one if and only if f = g.

4.6. Oplax normal double categories. Not much can be done with oplax morphisms that are not normal, and the more interesting construction of [12] was ℙath_{∗} for oplax normal double categories. This leads us to consider the notions of (strong) representability of composites and identities in oplax double categories. The definition of strongly representable composites in lax double categories was given in [12], Definition 2.7. We will give the oplax version here for convenience:

4.7. Definition. Let 𝔸 be an oplax double category and

[figure: A_{0}; f_{1}; A_{1}; f_{2}; f_{n}; A_{n}]

a path of horizontal arrows of 𝔸. We say that the composite of this path is strongly representable if there is an arrow f: A_{0} → A_{n} and a cell

[figure: A_{0}; f; η_{⟨fn,...,f_{1}⟩}; A_{n}; A_{0}; f_{1}; A_{1}; f_{2}; f_{n}; A_{n}]

such that for any paths ⟨x_{1}, …, x_{m}⟩, and ⟨y_{p}, …, y_{1}⟩, and cell α as below

[figure: B; g; •; v; α; C; •; w; X_{m}; x_{m}; X_{1}; x_{1}; A_{0}; f_{1}; A_{1}; f_{2}; f_{n}; A_{n}; y_{1}; Y_{1}; y_{2}; y_{p}; Y_{p}]

there is a unique cell \overline{α}

[figure: B; g; •; v; \overline{α}; C; •; w; X_{m}; x_{m}; X_{1}; x_{1}; A_{0}; f; A_{n}; y_{1}; Y_{1}; y_{2}; Y_{p}]

such that the composition of

[figure: B; g; •; v; \overline{α}; C; •; w; X_{m}; x_{m}; id; X_{m−1}; x_{m−1}; X_{1}; x_{1}; id; A_{0}; f; η_{⟨fm,...,f_{1}⟩}; A_{n}; y_{1}; id; Y_{1}; y_{2}; y_{p}; Y_{p}; X_{m}; x_{m}; X_{m−1}; x_{m−1}; X_{1}; x_{1}; A_{0}; f_{1}; A_{1}; f_{2}; f_{n}; A_{n}; y_{1}; Y_{1}; y_{2}; y_{p}; Y_{p}]

is equal to α.

We say that the composite of the path ⟨f_{n}, …, f_{1}⟩ is representable if we require the above condition only in the case where m = p = 0, i.e., when there are no x’s or y’s.

This definition was originally given by Lambek [25] for multicategories and by Hermida [21] for lax bicategories (there called multicategories with several objects) and by Leinster [26] for lax double categories (there called fc-multicategories). A different approach is given in Penon’s article [29] describing and extending Burroni’s work [5] on T -categories. Hermida’s work on representable multicategories [22] gives a more general notion of representability where the universal cell is not required to be special. However, since we want to think of the domain of the universal cell as the composite of its codomain sequence of composable arrows, we require the cells to be special. If vertical isomorphisms have both companions and conjoints, the two definitions are equivalent. It can also be verified that if all n-tuple composites are representable and the structure cells are closed under composition as required in [22], then all these composites are strongly representable.

When the composite of a string [figure: A_{0}; f_{1}; A_{1}; f_{n}; A_{n}] is strongly representable we will choose a representing morphism and call it f_{n}f_{n−1} ⋯ f_{1}. We will also say that the composite exists in this case. The representing cell is denoted

[figure: A_{0}; f_{n}···f_{1}; ι_{fn,f_{n−1},...,f_{1}}; A_{n}; A_{0}; f_{1}; A_{1}; f_{2}; f_{n}; A_{n}.]

For n = 0, the composite, when it exists, will be denoted by 1_{A} and we say that the identity on A exists or is strongly representable. The universal cell is denoted

[figure: A; 1_{A}; ι_{A}; A; A; .]

As is historically the case, we are mostly interested in identities and composites of two arrows.

The following proposition is due to Hermida [21], Theorem 11.6, although it is stated for oplax bicategories in that paper.

4.8. Proposition. 𝕊pan A has all composites if and only if A has pullbacks.

We shall need refinements of this result for categories which may not have all pullbacks.

4.9. Proposition. For any A, horizontal identities exist in 𝕊pan A. Moreover, for any functor F: A → B, the functor 𝕊pan (F) preserves these identities.

Proof. It is easily verified that the cell

[figure: A; A; 1_{A}; 1_{A}; A; A]

has the required universal property.

Following Definition 3.15 of [12] we say that an oplax double category is normal if all identities exist and an oplax morphism is normal if it preserves these identities in the sense that

[figure: F A; F 1_{A}; F ι_{A}; F A; F A]

has the universal property making F 1_{A} an identity on F A. Thus, 𝕊pan A and 𝕊pan (F) are normal.

4.10. Proposition. Let A ⟵^{s_{1}} S ⟶^{s_{2}} B and B ⟵^{t_{1}} T ⟶^{t_{2}} C be spans in A. Then the following are equivalent.

1. T S is representable;

2. T S is strongly representable;

3. The pullback \xymatrix{ & P \ar[dl]_{s} \ar[dr]^{t} \\ S \ar[dr]_{s_{2}} & & T \ar[dl]^{t_{1}} \\ & B} exists.

Furthermore, 𝕊pan (F): 𝕊pan A → 𝕊pan B preserves the composite T S if F preserves the pullback S ×_{B} T.

Proof. To prove the implication (1)⇒(3), suppose that T S is representable with universal cell ι as in the diagram

[figure: A; P; s; t; C; A; S; s_{1}; s_{2}; B; T; t_{1}; t_{2}; C; .]

For a commutative square \xymatrix{ & X \ar[dl]_{x} \ar[dr]^{y} \\ S \ar[dr]_{s_{2}} & & T \ar[dl]^{t_{1}} \\ & B} we get a cell

[figure: A; X; s_{1}x; t_{2}y; x; y; C; A; S; s_{1}; s_{2}; B; T; t_{1}; t_{2}; C]  (24)

which must factor uniquely through ι, i.e., there is a unique z such that

[figure: A; X; z; C; A; P; s; t; C; A; S; s_{1}; s_{2}; B; T; t_{1}; t_{2}; C]

is equal to (24). Thus, \xymatrix{ & P \ar[dl]_{s} \ar[dr]^{t} \\ S \ar[dr]_{s_{2}} & & T \ar[dl]^{t_{1}} \\ & B} is a pullback.

The implication (3)⇒(2) is immediate from the universal property of pullbacks, and the implication (2)⇒(1) is obvious.

Finally, that 𝕊pan F preserves the composite T S when F preserves the pullback in Condition (3) is also obvious.

4.11. Corollary.

1. If g: B → C is an arrow in A, then g_{∗}S is strongly representable in 𝕊pan A.

2. If f: B → A is an arrow in A, then T f^{∗} is strongly representable in 𝕊pan A.

3. For any functor F: A → B, the induced morphism 𝕊pan (F) preserves composites of the form g_{∗}S and T f^{∗}.

Proof. The arrow g_{∗} is the span B ⟵^{1_{B}} B ⟶^{g} C and f^{∗} is the span A ⟵^{f} B ⟶^{1_{B}} B. In either case, the pullback of Condition (3) in Proposition 4.10 is of the form

\xymatrix{X \ar[d]_{1_{X}} \ar[r] & B \ar[d]^{1_{B}} \\ X \ar[r] & B ,}

which always exists for trivial reasons and is preserved by any functor.

4.12. Corollary. The composites 1_{B}S and S1_{A} exist and are canonically isomorphic to S.

We conclude that for any category A, the oplax double category 𝕊pan A is normal and for any functor F, 𝕊pan (F) is normal. Moreover, 𝕊pan A has composites of the form f_{∗}S and T f^{∗} and 𝕊pan (F) preserves them. If we want to extend the universal properties of Theorems 2.8 and 3.15 to oplax bicategories and double categories, representability of such composites has to be considered. For this we need a notion of gregarious oplax double category. There is a very natural extension of the notion of being gregarious to oplax normal double categories in terms of companions and conjoints. However, in order to be able to talk about companions and conjoints in oplax normal double categories we first need to consider composition of cells in more detail.

4.13. Composition of cells. We begin with a couple of results on composition of horizontal arrows.

4.14. Proposition. Let 𝔸 be an oplax double category and \xymatrix{A \ar[r]^{f} & B \ar[r]^{g} & C \ar[r]^{h} & D} horizontal arrows in 𝔸.

1. If gf and hg are strongly representable, the following are equivalent:

(a) h(gf) is strongly representable;

(b) (hg)f is strongly representable;

(c) hgf is strongly representable.

Furthermore, if any one of these is strongly representable, then we have h(gf) ≅ (hg)f ≅ hgf by vertically invertible special cells.

2. If 1_{B} is strongly representable, then so are the composites 1_{B}f and g1_{B}. Moreover, 1_{B}f ≅ f and g1_{B} ≅ g.

3. If any oplax morphism preserves 1_{B} then it preserves the composites 1_{B}f and g1_{B}.

Proof. These are all more or less standard representability results. Let us first show (1a) ⇔ (1c). Assume that gf, gh, and h(gf) are strongly representable with universal cells

[figure: A; gf; ι_{g,f}; C; B; hg; ι_{h,g}; D; A; h(gf); ι_{h,gf}; C; A; f; B; g; C; ,; B; g; C; h; D; ,; A; gf; C; h; D.]

Define ι_{h,g,f} to be the composite

[figure: A; h(gf); ι_{h,gf}; D; A; gf; ι_{g,f}; C; h; _{id_{h}}; D; A; f; B; g; C; h; D.]

The universal property of ι_{h,g,f} is easily checked, but note that strong representability of gf is needed, even if we only want mere representability of hgf. So we see that h(gf) strongly represents hgf. By the universal property for representability, if we had chosen any other arrow hgf to represent this composite, we conclude that h(gf) ≅ hgf by a vertically invertible special cell.

Assume now that gf, hg, and hgf are strongly representable. Strong representability of gf tells us that there is a unique cell ι_{h,gf} such that

[figure: A; hgf; ι_{h,gf}; D; A; gf; ι_{g,f}; C; h; _{id_{h}}; D; =; A; hgf; ι_{h,g,f}; D; A; f; B; g; C; h; D; A; f; B; g; C; h; D.]

It is again straightforward to check that ι_{h,gf} has the universal property making h(gf) strongly represented by hgf, and h(gf) ≅ hgf. The rest of (1) follows by duality.

For (2), assume that 1_{B} is strongly representable with universal cell

[figure: B; 1_{B}; ι_{B}; B; B; .]

Then there is a unique cell ι_{g,1_{B}} such that

[figure: B; ι_{g,1_{B}}; g; C; B; _{id_{g}}; g; C; B; 1_{B}; ι_{B}; B; _{id_{g}}; g; C; =; B; g; C.; B; g; C]

It is easily seen that ι_{g,1_{B}} has the universal property which makes g1_{B} strongly represented by g. The proof of the fact that 1_{B}f is strongly representable is dual to this.

For (3), assume that F preserves 1_{B}. Then strongly represents the horizontal identity on F B. So there is a unique cell ι_{F g,F 1_{B}} such that

[figure: F B; F 1_{B}; F ι_{B}; F B; F B]

[figure: F B; ι_{F g,F 1_{B}}; F g; F C; F B; _{id_{F g}}; F g; F C; F B; F 1_{B}; F ι_{B}; F B; _{id_{F g}}; F g; F C; =; F B; F g; F C.; F B; F g; F C]

Just as in the proof of the previous part, it is easily seen that ι_{F g,F 1_{B}} has the universal property which makes F gF 1_{B} strongly represented by F g. The proof for the preservation of 1_{B}f is dual to this one.

Note that the above proof shows that h(gf) satisfies the universal property of hgf, so we could choose hgf to be h(gf). However, we cannot choose it to be (hg)f at the same time. It is best to make an independent choice of hgf. Then we get a unique isomorphism κ: h(gf) → hgf satisfying

[figure: A; h(gf); κ; D; A; ι_{h,gf}; h(gf); D; A; ι_{h,g,f}; hgf; D; =; A; ι_{g,f}; gf; C; h; D; _{id_{h}}; A; f; B; g; C; h; D; A; f; B; g; C; h; D.]

The horizontal composition of arrows in an oplax double category requires the solution of a universal mapping problem, and when it does exist associativity and the unit laws only hold up to comparison cells which are vertical isomorphisms if we have strong representability. However, we will see that horizontal composition of cells works perfectly well provided that the horizontal composites of the relevant domains and codomains exist. All reasonable laws hold on the nose once the canonical isomorphisms are factored in.

There are two ways we might define horizontal composition of cells in an oplax double category. The first is more general and a graded kind of composition.

Given cells α and β as below where gf is strongly representable, the horizontal composite β ∗ α is defined to be the vertical composite (β, α) · ι_{g,f},

[figure: A; •; α; f; B; •; β; g; C; •; X_{0}; X_{1}; X_{k}; X_{n},]

[figure: A; ι_{g,f}; gf; C; A; •; f; α; B; g; •; β; C; •; X_{0}; X_{k}; X_{n}.]

Similarly, given a vertical arrow \xymatrix{v: A \ar[r]|{•} & X}, where 1_{A} is strongly representable, the horizontal identity cell 1_{v}^{∗} is defined to be the composite

[figure: A; ι_{A}; A; A; •; v; X.]

4.15. Proposition.

1. Given cells α, and β as above and a further one

[figure: C; h; γ; •; D; •; X_{n}; X_{m},]

such that gf, hg, and hgf are strongly representable, then γ ∗ (β ∗ α) = (γ ∗ β) ∗ α, in the sense that

[figure: A; hgf; ∥≀; D; A; hgf; ∥≀; D; A; •; h(gf); γ∗(β∗α); D; •; =; A; •; (hg)f; (γ∗β)∗α; D; •; X_{0}; X_{m}; X_{0}; X_{m}]

where the top cells are the canonical isomorphisms mentioned in Proposition 4.14.

2. If 1_{B} is strongly representable, then 1_{v}^{∗} ∗ α = α and β ∗ 1_{v}^{∗} = β, modulo the canonical isomorphisms 1_{B}f ≅ f and g1_{B} ≅ g as in Part (1).

3. With cells α and β as above, and gf representable, there is also an interchange law.

Given cells

[figure: X_{i−1}; •; ξ_{i}; X_{i}; •; Y_{j_{i−1}}; Y_{j_{i}},]

we have

[(ξ_{n}, ξ_{n−1}, …, ξ_{k}) · β] ∗ [(ξ_{k}, …, ξ_{1}) · α] = (ξ_{n}, …, ξ_{1}) · (β ∗ α).

Proof. Straightforward. (The proof of Part (3) is easiest, being nothing but associativity of vertical composition in 𝔸.)

There is another more special kind of horizontal composition on cells with a square boundary, i.e., for which the vertical codomain has length 1. This notion of composition is the one which is more useful for us.

Given cells α and β as in

[figure: A; f; α; •; u; B; •; v; B; •; v; g; β; C; •; w; X; f′; Y; Y; g′; Z]

such that gf and g′f′ are strongly representable, the horizontal composite βα is the unique cell such that

[figure: A; gf; •; u; βα; C; •; w; A; gf; ι_{g,f}; C; X; g′f′; ι_{g′,f′}; Z; =; A; •; u; f; α; B; •; v; β; g; C; •; w; X; f′; Y; g′; Z; X; f′; Y; g′; Z,]  (25)

i.e., ι · (βα) = β ∗ α.

Also, given a vertical arrow \xymatrix{v: B \ar[r]|{•} & Y}, for which 1_{B} and 1_{Y} are both strongly representable, we define the horizontal identity cell 1_{v} to be the unique cell such that 4.16. Proposition.

[figure: B; •; v; 1_{B}; 1_{v}; B; •; v; B; 1_{B}; ι_{B}; B; Y; 1_{Y}; ι_{Y}; Y; =; B; •; v; Y; Y; .]

1. Given cells α and β as above, and a further one

[figure: C; •; w; h; γ; D; •; r; Z; h′; W]

for which gf, hg, hgf, g′f′, h′g′, and h′g′f′ are all strongly representable, then γ(βα) = (γβ)α (modulo the canonical isomorphisms).

2. If 1_{B} and 1_{Y} are strongly representable, then 1_{v}α = α and β1_{v} = β.

3. (The Interchange Law) If we also have cells

[figure: X; •; u′; f′; α′; Y; •; v′; and; Y; •; v′; g′; β′; Z; •; w′; A′; f′′; B′; B′; g′′; C′,]

and gf, g′f′, and g′′f′′ are all strongly representable, then

(β′α′) · (βα) = (β′ · β)(α′ · α).

Proof. There is a triple horizontal composite of cells γβα defined to be the unique cell such that

[figure: A; •; γβα; hgf; D; •; A; ι_{f,g,h}; hgf; D; X; h′g′f′; ι_{f′,g′,h′}; W; =; A; •; α; f; B; •; β; g; C; •; γ; h; D; •; X; f′; Y; g′; Z; h′; W; X; f′; Y; g′; Z; h′; W.]

We will show that γ(βα) = γβα in the sense that

[figure: A; •; γ(βα); h(gf); D; •; A; κ; h(gf); D; X; κ; h′(g′f′); W; =; A; hgf; •; γβα; D; •; X; h′g′f′; W; X; h′g′f′; W.]

Then the result will follow by duality. These cells will be equal if they are equal when composed (vertically) with ι_{h′,g′,f′}. Consider the following diagrams whose composites are all equal.

[figure: A; •; γ(βα); h(gf); D; •; A; •; γ(βα); h(gf); D; •; X; κ; h′(g′f′); W; =; X; h′(g′f′); ι_{h′,g′f′}; W; X; h′g′f′; ι_{h′,g′,f′}; W; X; ι_{g′,f′}; g′f′; Z; h′; W; _{id_{h′}}; X; f′; Y; g′; Z; h′; W; X; f′; Y; g′; Z; h′; W; A; h(gf); ι_{h,gf}; D; A; ι_{h,gf}; h(gf); D; =; A; •; gf; αβ; C; •; γ; h; D; •; =; A; gf; ι_{g,f}; C; h; D; _{id_{h}}; X; g′f′; ι_{g′,f′}; Z; h′; W; _{id_{h′}}; A; •; f; α; B; •; g; β; C; •; h; γ; D; •; X; f′; Y; g′; Z; h′; W; X; f′; Y; g′; Z; h′; W; A; h(gf); κ; D; A; h(gf); κ; D; =; A; hgf; ι_{h,g,f}; D; =; A; •; γβα; hgf; D; •; A; •; f; α; B; •; β; g; C; •; γ; h; D; •; X; h′g′f′; ι_{h′,g′,f′}; W; X; f′; Y; g′; Z; h′; W; X; f′; Y; g′; Z; h′; W.]

This proves Part (1).

The proof of Part (2) is similar and left as an exercise.

We give the proof of Part (3) which is a bit different. The cells (β′α′) · (βα) and (β′ · β)(α′ · α) are equal if they are equal when composed with ι_{g′,f′}. This is indeed the case, as the following sequence of diagrams with equal composites shows.

[figure: A; •; (β′α′)·(βα); C; •; A; •; βα; C; •; A; •; βα; C; •; A′; ι; C′; =; X; •; β′α′; Z; •; =; X; ι; Z; =; A′; B′; C′; A′; B′; C′; X; •; α′; Y; •; β′; Z; •; A′; B′; C′]

[figure: A; ι; C; A; ι; C; A; •; (β′·β)(α′·α); C; •; A; •; α; B; •; β; C; •; =; A; •; α′·α; B; •; β′·β; C; •; =; A′; ι; C′; X; •; α′; Y; •; β′; Z; •; A′; B′; C′; A′; B′; C′.; A′; B′; C′]

4.17. Corollary. If in any oplax double category all composites and all identities are strongly representable, the square cells (with vertical codomains of length one) form a (weak) double category.

4.18. Gregarious oplax double categories. For an oplax normal double category the definitions of companions and conjoints given in Definition 3.4 for double categories still make sense:

4.19. Definition. Let 𝔹 be an oplax normal double category. We say that a horizontal arrow \xymatrix{f: A \ar[r] & B} and a vertical arrow \xymatrix{v: A \ar[r]|{•} & B} in 𝔹 are companions if there are binding cells

[figure: A; I_{A}; ψ; A; •; v; and; A; •; v; f; χ; B; A; f; B; B; I_{B}; B,]

such that

[figure: A; I_{A}; ψ; A; •; v; A; I_{A}; ψ; A; •; v; χ; f; B; = id_{f}; A; •; v; χ; f; B; = 1_{v}; and; A; f; B; I_{B}; B; B; I_{B}; B; ,]  (26)

i.e., χ · ψ = 1_{v} and χψ = id_{f}.

Dually, we say that a horizontal arrow \xymatrix{u: B \ar[r] & A} and a vertical arrow \xymatrix{v: A \ar[r]|{•} & B} are conjoints if there are binding cells

[figure: A; •; v; α; I_{A}; A; and; B; u; β; A; •; v; B; u; A; B; I_{B}; B,]

such that

[figure: A; •; v; I_{A}; α; A; B; u; β; A; •; v; α; I_{A}; A; = id_{f}; B; β; u; A; •; v; = 1_{v}; and; B; I_{B}; B; u; A; B; I_{B}; B; ,]

i.e., β · α = 1_{v} and αβ = id_{f}.

Note that these definitions only use composites of two horizontal arrows in which one is an identity and these exist by Proposition 4.14 Part (2). They also use horizontal composition of cells and horizontal identities which exist and have the required properties by Proposition 4.16. Note however, that the definition of adjoint given in Definition 3.4 does not carry over to the present context as it uses the triple composites fuf and ufu.

To illustrate how the special horizontal composition of cells works here, we spell out what is meant by the equations in (26). The left-hand equation means that

[figure: I; ψ; •; v; I; ι; •; v; f; χ; =; •; v; I; ι]

whereas the right-hand equation means that

[figure: f; ι_{f,I}; f; _{id_{f}}; ψ; I; •; v; χ; f; =; f; ι_{I,f}; f; I; f; I]

We derive from this that

[figure: f; ι_{f,I}; I; ψ; •; v; χ; f; =; _{id_{f}}; f; f; I; ι; f; f]  (27)

The following properties of conjoints and companions hold in any oplax normal double category.

1. If h and h′ are both companions to v then h ≅ h′ by a special vertical isomorphism.

2. If h is companion to v and v′ then v ≅ v′ by a special horizontal isomorphism.

3. If \xymatrix{A \ar[r]|{•}^{v} & B \ar[r]|{•}^{w} & C} are vertical morphisms with companions h and k respectively, then w · v has a companion if and only if the composite kh is strongly representable and that composite is the companion of w · v.

4. The vertical arrow id_{A} has the horizontal arrow 1_{A} as companion.

5. Any morphism of oplax normal double categories preserves companions.

6. Dual statements hold for conjoints. We are now ready to introduce gregarious oplax double categories and discuss the

universal properties of 𝕊pan A for an arbitrary category A. 4.20. Definition. An oplax double category is gregarious if it satisfies the following conditions.

1. It is normal.

2. Every vertical arrow v has a horizontal companion v_{∗} and a horizontal conjoint v^{∗}.

3. For each span of vertical arrows \xymatrix{A & B \ar[l]_{v} \ar[r]^{w} & C}, the composite w_{∗}v^{∗} is strongly representable.

We write GregOplax for the 2-category of gregarious oplax double categories, normal morphisms and vertical transformations.

4.21. Example. For any category A, the oplax double category 𝕊pan A is gregarious. Also, for any functor F: A → B, the induced functor 𝕊pan (F) of oplax double categories is a normal morphism, as was shown above.

The morphisms between gregarious oplax double categories are not explicitly required to preserve the composites of the form w_{∗}v^{∗}, because normal morphisms will do this automatically.

4.22. Proposition. A normal morphism F: ℂ → 𝔻 between oplax normal double categories preserves all composites of the forms w_{∗}x and yw^{∗} that exist in ℂ.

Proof. We will show that F preserves the composites of the form w_{∗}x that exist in ℂ. The dual of this proof shows that F also preserves the composites of the form yw^{∗} that exist in ℂ. Let

[figure: w_{∗}x; ι_{w∗,x}; x; w_{∗}]

be a universal cell with the strong representability property in ℂ for the composite w_{∗}x. We will now show that

[figure: F (w_{∗}x); F ι_{w∗,x}; F x; F w_{∗}]

has the strong representability property in 𝔻.

Let

[figure: g; β; a_{1}; ⋯; a_{n}; F x; F w_{∗}; b_{1}; ⋯; b_{m}]

be any cell in 𝔻. We need to show that there is a unique cell β̄ such that β is the composite,

[figure: g; β̄; a_{1}; ⋯; a_{n}; F ι_{w∗,x}; F (w_{∗}x); b_{1}; ⋯; b_{m}; a_{1}; ⋯; a_{n}; F x; F w_{∗}; b_{1}; ⋯; b_{m}; .]  (28)

362  ROBERT DAWSON, ROBERT PARÉ, DORETTE PRONK

For the sake of space, we will from now on abbreviate the path [figure: a_{1}; ⋯; a_{n}]by [figure: (a_{i})] (and analogously for the (b_{j}) path.) We define β̄ to be the composite of

[figure: g; β; (a_{i}); F (ψ_{w}id_{x}); F x; •; F w; F χ_{w}; F w_{∗}; (b_{j}); (a_{i}); F (w_{∗}x); F I; F ι; (b_{j}); (a_{i}); F (w_{∗}x); (b_{j}); .]

Recall that the cell ψ_{w}id_{x} is defined as the unique cell such that

[figure: x; ι_{I,x}; x; ψ_{w}id_{x}; •; w; _{id_{x}}; x; ψ_{w}; I; •; w; =; w_{∗}x; ι_{w∗,x}; x; w_{∗}; x; w_{∗}; .]  (29)

We will now calculate the composite (28) and show that it is indeed equal to β.

By the associativity and identity properties of the composition of cells in an oplax double category, this is equal to By (29) this is equal to

[figure: g; β̄; g; β; (a_{i}); F ι_{w∗,x}; F (w_{∗}x); (b_{j}); =; (a_{i}); F x; F (ψ_{w}id_{x}); •; F w; F w_{∗}; F χ_{w}; (b_{j}); (a_{i}); F x; F w_{∗}; (b_{j}); (a_{i}); F (w_{∗}x); id; F I; F ι; (b_{j}); (a_{i}); F (w_{∗}x); F ι_{w∗,x}; (b_{j}); (a_{i}); F x; F w_{∗}; (b_{j}); .]

[figure: g; β; (a_{i}); F x; F (ψ_{w}id_{x}); •; F w; F χ_{w}; F w_{∗}; (b_{j}); (a_{i}); F (w_{∗}x); F ι_{w∗,x}; F I; F ι; (b_{j}); (a_{i}); F x; F w_{∗}; (b_{j}); .]

[figure: g; β; (a_{i}); F x; F ι_{I,x}; F w_{∗}; (b_{j}); (a_{i}); F x; F I; F ψ_{w}; •; F w; F χ_{w}; F w_{∗}; (b_{j}); (a_{i}); F x; F w_{∗}; F I; F ι; (b_{j}); (a_{i}); F x; F w_{∗}; (b_{j}); .]

By the universal property of

[figure: F I; F ι; ,]

this is equal to

[figure: g; β; (a_{i}); F x; F w_{∗}; F ι_{w∗,I}; (b_{j}); (a_{i}); F x; F I; F ψ_{w}; •; F w; F χ_{w}; F w_{∗}; (b_{j}); (a_{i}); F x; F w_{∗}; F I; F ι; (b_{j}); (a_{i}); F x; F w_{∗}; (b_{j}); .]

By (27) this is equal to β.

It remains to be shown that β̄ is unique with this property. So suppose that γ is a cell satisfying

[figure: g; γ; =; g; β; (a_{i}); F (w_{∗}x); F ι_{w∗,x}; (b_{j}); (a_{i}); F x; F w_{∗}; (b_{j}); .; (a_{i}); F x; F w_{∗}; (b_{j})]  (30)

This implies that β̄ is equal to the composite

[figure: g; γ; (a_{i}); F (w_{∗}x); F ι_{w∗,x}; (b_{j}); (a_{i}); F x; F (ψ_{v}id_{x}); •; F w; F χ_{w}; F w_{∗}; (b_{j}); (a_{i}); F (w_{∗}x); F I; F ι; (b_{j}); (a_{i}); F (w_{∗}x); (b_{j}); .]  (31)

Now consider the following composite in ℂ:

[figure: w_{∗}x; ι_{w∗,x}; x; ι_{I,x}; w_{∗}; x; I; ψ_{w}; •; w; χ_{w}; w_{∗}; x; w_{∗}; I; ι; x; w_{∗}; .]  (32)

By coherence and (27) this is equal to By (29) the composite (32) is also equal to

[figure: w_{∗}x; ι_{w∗,x}; x; w_{∗}; ι_{w∗,I}; =; w_{∗}x; ι_{w∗,x}; x; I; ψ_{w}; •; w; w_{∗}; χ_{w}; x; w_{∗}; x; w_{∗}; I; x; w_{∗}; .]

[figure: w_{∗}x; ι_{w∗,x}; w_{∗}x; ι_{w∗,x}; x; ψ_{w}id_{x}; •; v; χ_{w}; w_{∗}; =; x; ψ_{w}id_{x}; •; v; χ_{w}; w_{∗}; w_{∗}x; ι_{w∗,x}; I; ι; w_{∗}x; I; ι; x; w_{∗}; w_{∗}x; ι_{w∗,x}; x; w_{∗}; .]

By the universal property of

[figure: w_{∗}x; ι_{w∗,x}; x; w_{∗}]

this implies that

[figure: w_{∗}x; ι_{w∗,x}; x; ψ_{w}id_{x}; •; v; χ_{w}; w_{∗}; =; w_{∗}x; id; w_{∗}x; I; ι; w_{∗}x; w_{∗}x; .]

Substituting the image under F of this into (31), we get that that composite is equal to γ. So we conclude that β̄ = γ as desired.

4.23. Corollary. Any normal morphism between gregarious oplax double categories preserves all composites of the form w_{∗}v^{∗} for vertical arrows w and v with a common domain.

4.24. The Main Theorem. We will now show that 𝕊pan A is the universal gregarious oplax double category in the following sense.

4.25. Theorem. Let 𝔹 be a gregarious oplax double category. Composition with the vertical inclusion A ↪ 𝕊pan A induces an equivalence of categories between the category of normal morphisms 𝕊pan A → 𝔹 and (vertical) transformations and the category of functors A → Vrt 𝔹 and natural transformations,

GregOplax(𝕊pan A, 𝔹) ≃ Cat (A, Vrt 𝔹).

Before we prove this theorem we want to introduce the canonical cells associated with a commutative square of vertical arrows

[figure: x; u; v; y]

in a gregarious oplax double category. We first reintroduce some of the cells from (9). Note that both ζ(x, u, v, y) and ξ(x, u, v, y) can be defined using the special horizontal composition of square cells defined in (25):

ζ(x, v, u, y) = (α_{y} · 1_{u})(1_{v} · β_{x}) and ξ(x, v, u, y) = (1_{y} · χ_{u})(ψ_{v} · 1_{x}).

In addition to these cells, we will need the cells

[figure: ι; •; x; •; x; •; u; •; u; Λ(x, u, v, y); :=; ψ_{v}; •; v; •; y; α_{y}; v_{∗}; y^{∗}; .]

Note that for any commutative diagram of vertical arrows

[figure: u_{1}; x; u_{2}; v_{1}; y; v_{2}; z]

we have that

ζ(y, v_{1}, v_{2}, z) · ζ(x, u_{1}, u_{2}, y) = ζ(x, v_{1} · u_{1}, v_{2} · u_{2}, z),  (33)

and

ξ(y, v_{1}, v_{2}, z) · ξ(x, u_{1}, u_{2}, y) = ξ(x, v_{1} · u_{1}, v_{2} · u_{2}, z).  (34)

Also, for a commutative diagram of vertical arrows

[figure: u; x; v; r; w; t; y; s; z]

it is straightforward to check that

(ζ(y, t, s, z), ξ(v, s, r, w)) · Λ(x, u, v, y) = Λ(r · x, t · u, w, z).  (35)

Proof. (of Theorem 4.25) Since 𝔹 is gregarious, start by choosing companions and conjoints with their binding cells for all vertical arrows. For vertical identities choose (id_{A})^{∗} = 1_{A} = (id_{A})_{∗} with ι_{A} as binding cells. Also choose the composites w_{∗}v^{∗} for all spans of vertical arrows.

To show that the functor defined by composition is essentially surjective, let F: A → Vrt 𝔹 be any functor. We will now construct its extension to a normal morphism of oplax double categories F̃: 𝕊pan A → 𝔹. Obviously, on objects and vertical arrows F̃ has the same values as F.

On horizontal arrows, F̃^{( [figure: q; p] )} = (F p)_{∗}(F q)^{∗} (the chosen representative for this strongly representable horizontal composition).

The image F̃(μ) of a cell

[figure: μ =; A; u; S; q; p; v_{1}; v_{2}; v_{m}; B; w; C_{0}; S_{1}; q_{1}; p_{1}; C_{1}; S_{2}; q_{2}; p_{2}; C_{2}; S_{m}; q_{m}; p_{m}; C_{m}]  (36)

under F̃ is defined as

[figure: (F p)_{∗}(F q)^{∗}; ι; •; F u; (F q)^{∗}; ζ; •; F v_{1}; Λ; •; F v_{2}; •; F v_{m−1}; Λ; •; F v_{m}; (F p)_{∗}; ξ; •; F w; (F q_{1})^{∗}; (F p_{1})_{∗}; (F q_{2})^{∗}; (F p_{m})_{∗}; (F q_{m})^{∗}; (F p_{m})_{∗}; .]

It is straightforward to check that F̃ is normal.

It is obvious that composition with the inclusion of A into 𝕊pan A defines a faithful functor GregOplax(𝕊pan A, 𝔹) → Cat (A, Vrt 𝔹). To show that it is also full, let γ: F → G be a natural transformation. We will now construct its extension γ̃: F̃ → G̃. For objects A in 𝕊pan A, γ̃_{A} = γ_{A}, taken as a vertical arrow. For a span \xymatrix{A & S \ar[l]_{q} \ar[r]^{p} & B}, define

γ̃_{(q,S,p)} = ξ(F p, γ_{S}, γ_{B}, Gp)ζ(F q, γ_{S}, γ_{A}, Gq),

i.e., the special horizontal composition of

[figure: F A; •; γ_{A}; ζ; (F q)^{∗}; F S; •; γ_{S}; (F p)_{∗}; ξ; F B; •; γ_{B}; GA; (Gq)^{∗}; GS; (Gp)_{∗}; GB .]

To verify vertical naturality of γ̃, we need to show that

G̃(μ) · γ̃_{(q,S,p)} = (γ̃_{(q_{m},S_{m},p_{m})}, ⋯, γ̃_{(q_{1},S_{1}.p_{1})}) · F̃(μ)  (37)

for any cell μ as in (36). The left-hand side of this equation is the following cell:

[figure: •; γ̃_{A}; (F p)_{∗}(F q)^{∗}; ξζ; •; γ̃_{B}; (Gp)_{∗}(Gq)^{∗}; ι; •; Gu; (Gq)^{∗}; ζ; •; Gv_{1}; Λ; •; Gv_{2}; •; Gv_{m−1}; Λ; •; Gv_{m}; ξ; (Gp)_{∗}; •; Gw; (Gq_{1})^{∗}; (Gp_{1})_{∗}; (Gq_{2})^{∗}; (Gp_{m−1})_{∗}; (Gq_{m})^{∗}; (Gp_{m})_{∗}; .]

By the definition of special horizontal composition in (25), this can be rewritten as

[figure: (F p)_{∗}(F q)^{∗}; ι; •; γ̃_{A}; ζ; (F q)^{∗}; •; γ̃_{S}; •; γ̃_{S}; •; γ̃_{S}; •; γ̃_{S}; ξ; (F p)_{∗}; •; γ̃_{B}; •; Gu; ζ; (Gq)^{∗}; •; Gv_{1}; Λ; •; Gv_{2}; •; Gv_{m−1}; Λ; •; Gv_{m}; ξ; (Gp)_{∗}; •; Gw; (Gq_{1})^{∗}; (Gp_{1})_{∗}; (Gq_{2})^{∗}; (Gp_{m−1})_{∗}; (Gq_{m})^{∗}; (Gp_{m})_{∗}; .]  (38)

Note that

Λ(Gv_{i}, Gv_{i+1}, Gp_{i}, Gq_{i+1}) · 1γ̃_{S} = Λ(Gv_{i} · γ̃_{S}, Gv_{i+1} · γ̃_{S}, Gp_{i}, Gq_{i+1}).

Using this result, together with (33) and (34), we can rewrite (38) as

[figure: (F p)_{∗}(F q)^{∗}; ι; •; γ̃_{A}; ζ; (F q)^{∗}; •; γ̃_{S}; Λ; •; γ̃_{S}; •; γ̃_{S}; Λ; •; γ̃_{S}; ξ; (F p)_{∗}; •; γ̃_{B}; •; Gu; •; Gv_{1}; •; Gv_{2}; •; Gv_{m−1}; •; Gv_{m}; •; Gw; (Gq_{1})^{∗}; (Gp_{1})_{∗}; (Gq_{2})^{∗}; (Gp_{m−1})_{∗}; (Gq_{m})^{∗}; (Gp_{m})_{∗}; .]

Since Gu · γ̃_{A} = γ̃_{C_{0}} · F U, Gw · γ̃_{B} = γ̃_{C_{m}} · F w, and

Λ(Gv_{i} · γ̃_{S}, Gv_{i+1} · γ̃_{S}, Gp_{i}, Gq_{i+1}) = Λ(γ̃_{S_{i}} · F v_{i}, γ̃_{S} · F v_{i+1}, Gp_{i}, Gq_{i+1}) we can further rewrite this as

[figure: (F p)_{∗}(F q)^{∗}; ι; •; F u; ζ; (F q)^{∗}; •; F v_{1}; Λ; •; F v_{2}; •; F v_{m−1}; Λ; •; F v_{m}; ξ; (F p)_{∗}; •; F w; •; γ̃_{C_{0}}; •; γ̃_{S_{1}}; •; γ̃_{S_{2}}; •; γ̃_{S_{m−1}}; •; γ̃_{Sm}; •; γ̃_{Cm}; (Gq_{1})^{∗}; (Gp_{1})_{∗}; (Gq_{2})^{∗}; (Gp_{m−1})_{∗}; (Gq_{m})^{∗}; (Gp_{m})_{∗}; .]  (39)

Now we use (35) for the commutative diagram

[figure: •; F v_{i+1}; •; F v_{i}; •; F p_{i}; •; γ̃_{S_{i}}; •; Gp_{i}; •; γ̃_{S_{i+1}}; •; F q_{i+1}; •; γ̃_{C_{i}}; •; Gq_{i+1}]

together with (33) and (34) to rewrite (39) as

[figure: (F p)_{∗}(F q)^{∗}; ι; •; F u; ζ; (F q)^{∗}; •; F v_{1}; Λ; •; F v_{2}; •; F v_{m−1}; Λ; •; F v_{m}; ξ; (F p)_{∗}; •; F w; •; γ̃_{C_{0}}; (F q_{1})^{∗}; ζ; •; γ̃_{S_{1}}; (F p_{1})_{∗}; ξ; •; γ̃_{C_{1}}; (F q_{2})^{∗}; ζ; •; γ̃_{S_{2}}; •; γ̃_{S_{m−1}}; (F p_{m−1})_{∗}; ξ; •; γ̃_{C_{m−1}}; ζ; (F q_{m})^{∗}; •; γ̃_{Sm}; (F p_{m})_{∗}; ξ; •; γ̃_{Cm}; (Gq_{1})^{∗}; (Gp_{1})_{∗}; (Gq_{2})^{∗}; (Gp_{m−1})_{∗}; (Gq_{m})^{∗}; (Gp_{m})_{∗}; ,]

and this is the cell on the right-hand side of (37). So γ̃ forms a vertical natural transformation from F̃ to G̃. This concludes the proof of Theorem 4.25.

4.26. Corollary. 𝕊pan: Cat → GregOplax is locally fully faithful.

4.27. Observation. Since the embedding GregDoub_{OplaxN} ↪ GregOplax is 2-full and faithful, we can also view Theorem 3.15 as a corollary of Theorem 4.25. This gives us an alternative proof for Theorem 3.15, although at the cost of introducing additional concepts which are not needed in the context of that proof.

4.28. Free Adjoints Revisited. In this section we combine our 𝕊pan construction with the ℙath_{∗} construction of [12] to get a double category version of the Π_{2} construction of [14]. It will be important to distinguish between strict double categories and weak double categories and between strict and strong morphisms of double categories. The universal properties of ℙath and ℙath_{∗} are with respect to strict double categories and strict morphisms between them, whereas the universal property of 𝕊pan is of necessity the weak version. To combine the two constructions they should be of the same type. However, the strong version does not follow from the weak one and neither does the weak property automatically follow from the strong one.

In order to get the weak property from the strong one, we will use two strictification results. The first is a variation on the coherence theorem for bicategories [27]. Let 𝔹 be a weak double category and define 𝔹̄, the strictification of 𝔹, to have the same objects and vertical arrows as 𝔹, but with paths of horizontal arrows in 𝔹 as its horizontal arrows. A cell

\xymatrix{B_{0} \ar[d]|{•}_{x} \ar[r]_{α}^{b_{1}} & B_{1} \ar[r]^{b_{2}} & B_{2} \ar[r]^{b_{3}} & ⋯ \ar[r]^{b_{n}} & B_{n} \ar[d]|{•}^{y} \\ C_{0} \ar[r]_{c_{1}} & C_{1} \ar[r]_{c_{2}} & C_{2} \ar[r]_{c_{3}} & ⋯ \ar[r]_{c_{m}} & C_{m}}

is a cell in 𝔹 of the form

\xymatrix{B_{0} \ar[d]|{•}_{x} \ar[r]_{α}^{∏ b_{i}} & B_{n} \ar[d]|{•}^{y} \\ C_{0} \ar[r]_{∏ c_{j}} & C_{m} ,}

where

∏_{i=1}^{n} b_{i} = \{ \begin{matrix}1_{B_{0}} & if & n = 0 \\ b_{1} & if & n = 1 \\ b_{n} (∏_{i=1}^{n−1} b_{i}) & if & n > 1. \end{matrix}

So ∏_{i=1}^{n} b_{i} = b_{n}(⋯ (b_{3}(b_{2}b_{1}) ⋯), i.e., all brackets are at the right. Vertical composition of cells in 𝔹̄ is the same as in 𝔹. Horizontal compostion of paths is by concatenation, whereas horizontal composition of cells uses the associativity isomorphisms of 𝔹. The coherence theorem for bicategories says that there are canonical isomorphisms (special cells)

φ_{m,n}: ( ∏_{i=n+1}^{n+m} b_{i}) (∏_{i=1}^{n} b_{i}) → ∏_{i=1}^{n+m} b_{i}

which satisfy all reasonable commutation properties, in particular the obvious associativity and unit laws. Then the horizontal composition of 𝔹̄ is given by

[figure: B_{0}; φ_{m,n}^{−1}; ∏ b_{i}; B_{m+n}; B_{0}; •; α; ∏ b_{i}; B_{n}; •; α′; ∏ b_{i}; B_{m+n}; •; C_{0}; ∏ c_{j}; φ_{p,q}; C_{p}; ∏ c_{j}; C_{p+q}; C_{0}; ∏ c_{j}; C_{p+q} .]

4.29. Theorem. With these operations, 𝔹̄ is a strict double category which is equivalent to 𝔹 in the 2-category psDoub of pseudo double categories, pseudo functors and vertical transformations.

Proof. Vertical composition, being the same as in 𝔹, gives no problem. Interchange is easy - the φ_{p,q} and φ_{p,q}^{−1} in the middle cancel. Horizontal composition of arrows, being concatenation, is strictly associative. The only thing to check are the associativity and unit laws for horizontal composition of cells. Associativity is represented schematically by

[figure: φ^{−1}; φ^{−1}; φ^{−1}; id; id; φ^{−1}; α′′(α′α) =; •; α; •; α′; •; α′′; •; =; •; α; •; α′; •; α′′; •; = (α′′α′)α.; φ; id; id; φ; φ; φ]

The unit law from 𝔹 says that

[figure: •; 1; 1; •; α; •; 1; φ; 1; φ; =; •; α; •]

which implies that

[figure: φ^{−1}; •; 1; 1; •; α; •; =; •; α; •; 1; φ; .]

There are pseudo functors Φ: 𝔹 → 𝔹̄ and Ψ: 𝔹̄ → 𝔹 which are the identity on objects

and vertical arrows. For a horizontal morphism f: B → B′ in 𝔹, Φ(f) = ( \xymatrix{B \ar[r]^{f} & B′} )

the path of length 1. Φ on cells is the identity. Ψ of a path \xymatrix{B_{0} \ar[r]^{f_{1}} & B_{1} \ar[r]^{f_{2}} & ⋯ \ar[r]^{f_{n}} & B_{n}} is ∏ f_{i}: B_{0} → B_{n} and Ψ of a cell is the same cell in 𝔹 now considered as a cell with vertical domain and codomain ∏ f_{i} and ∏ g_{j} rather than the paths ⟨f_{i}⟩ and ⟨g_{j}⟩.

That Φ and Ψ are pseudo functors is easy to verify. The isomorphisms Φ(f′f) → Φ(f′)Φ(f) and Φ(1_{B}) → 1_{ΦB} are special cells in 𝔹̄ coming from identities in 𝔹, whereas the isomorphisms Ψ⟨f_{i}⟩ → Ψ⟨f′_{j}⟩ → Ψ(⟨f_{i}⟩⟨f′_{j}⟩) are the φ_{p,q} above and 1_{B} → Ψ⟨ ⟩ is the identity.

The composite ΨΦ is obviously the identity on 𝔹 but the isomorphism ΦΨ ≅ 1𝔹̄ can use some elaboration. This isomorphism is given by the vertical transformations t: \xymatrix{ΦΨ \ar[r] & 1𝔹̄} and u: \xymatrix{1𝔹̄ \ar[r] & ΦΨ}, defined by t(B) = id_{B}: \xymatrix{B \ar[r]|{•} & B} and t(⟨f_{i}⟩) is

[figure: B_{0}; ∏ f_{i}; _{id_{∏ fi}}; B_{n}; B_{0}; f_{1}; B_{1}; f_{2}; ⋯; f_{n}; B_{n} .]

It is straightforward to check that t is a vertical transformation with inverse u defined analogously.

4.30. Corollary. For any pseudo double category 𝔸, Φ and Ψ induce an equivalence of categories

psDoub(𝔸, 𝔹) ≃ psDoub(𝔸, 𝔹̄).

4.31. Remarks.

1. 𝔹̄ is the object part of a 2-functor (¯ ): psDoub → stDoub which is a 2-left adjoint to the inclusion stDoub → psDoub. It is not, however, a 2-equivalence or even a biequivalence as one might be tempted to believe. This is because for strict 𝔹, although Φ: 𝔹 → 𝔹̄ is an equivalence, it is not an equivalence in stDoub. On the other hand, we do get a biequivalence between psDoub and its full sub 2-category determined by the strict double categories.

2. If 𝔹 is gregarious then so is 𝔹̄.

ℙath 𝔸, ℙath_{∗}𝔸, and 𝔹̄ all share the property that they are strict double categories whose horizontal categories are free on a graph. This is at the heart of our next strictification result.

4.32. Proposition. Let 𝔸 and 𝔹 be strict double categories and assume that the category Hor(𝔸) is free on a graph. Then the inclusion

stDoub(𝔸, 𝔹) ↪ psDoub(𝔸, 𝔹)

is an equivalence of categories.

Proof. Let F: 𝔸 → 𝔹 be a pseudo double functor. Define \overline{F}: 𝔸 → 𝔹 to be the same as F on objects and vertical arrows. A horizontal morphism in 𝔸 is represented by a unique path [figure: A_{0}; a_{1}; A_{1}; a_{2}; a_{n}; A_{n}] of edges in the generating graph. Define \overline{F} (⟨a_{i}⟩) to be the composite F (a_{n}) ⋯ F (a_{1}) in 𝔹. This makes \overline{F} strict at the level of horizontal arrows. For a cell

[figure: A_{0}; α; •; a_{1}; A_{1}; a_{2}; a_{n}; A_{n}; •; A′_{0}; a′_{1}; A′_{1}; a′_{2}; a′_{m}; A′_{m} ,]

define \overline{F} (α) to be

[figure: F A_{0}; F a_{1}; φ; F A_{1}; F a_{2}; F a_{n}; F A_{n}; F A_{0}; •; F α; F (a_{n}···a_{1}); F A_{n}; •; F A′_{0}; φ^{−1}; F (a′_{m}···a′_{1}); F A′_{m}; F A′_{0}; F a′_{1}; F A′_{1}; F a′_{2}; F a′_{m}; F A′_{m}]

where φ is the structural isomorphism.

It is straightforward to check that vertical transformations from F to G are in bijective correspondence with vertical transformations between \overline{F} and \overline{G}, since they are completely determined by their components for the edges of the generating graph.

4.33. Remark. In [16] such double categories were shown to be the flexible algebras for a 2-monad on Cat(Graph), of internal categories in the category Graph of non-reflexive directed graphs, whose strict algebras are (strict) double categories. The result above can also be seen as fitting in with known results about flexible algebras: strict morphisms from a flexible algebra A to a strict algebra B are in bijective correspondence with pseudo morphisms from A to B (see also [3]).

Recall from [12] that the universal properties of ℙath and ℙath_{∗} are given by 2adjunctions. Involved in these adjunctions are Oplax, the 2-category of oplax double categories, oplax morphisms and vertical transformations, and OplaxN, its locally full sub 2-category of normal oplax double categories and normal morphisms. Then ℙath: Oplax → stDoub is left 2-adjoint to the ‘inclusion’ 𝕆plax: stDoub → Oplax and ℙath_{∗}: OplaxN → stDoub is left 2-adjoint to the ‘inclusion’ 𝕆plax_{∗}: stDoub → OplaxN. This means that if 𝔸 is oplax and 𝔹 strict we have an isomorphism of categories

stDoub(ℙath𝔸, 𝔹) ⟶^{≅} Oplax(𝔸, 𝔹),

and if 𝔸 is oplax normal, we have an isomorphism of categories,

stDoub(ℙath_{∗}𝔸, 𝔹) ⟶^{≅} OplaxN(𝔸, 𝔹).

The ‘inclusions’ 𝕆plax and 𝕆plaxN extend to psDoub in the obvious way.

4.34. Theorem. The functors ℙath: Oplax → psDoub and ℙath_{∗}: OplaxN → psDoub satisfy the following universal properties. If 𝔸 is oplax and 𝔹 is a pseudo double category, then the morphism 𝔸 → ℙath𝔸 induces, by composition, an equivalence of categories

psDoub(ℙath𝔸, 𝔹) ⟶^{≃} Oplax(𝔸, 𝔹)

and if 𝔸 is oplax normal, composition with 𝔸 → ℙath_{∗}𝔸 induces an equivalence of categories

psDoub(ℙath_{∗}𝔸, 𝔹) ⟶^{≃} OplaxN(𝔸, 𝔹)

Proof. The proofs are similar, so we only do the case of ℙath_{∗} as it is the one we will use. We get the following equivalences

psDoub(ℙath_{∗}𝔸, 𝔹) ⟶^{≃} psDoub(ℙath_{∗}𝔸, \overline{𝔹})  by Corollary 4.30
⟶^{≃} stDoub(ℙath_{∗}𝔸, \overline{𝔹}) by the univ. prop. of ℙath_{∗}
⟶^{≃} OplaxN(𝔸, \overline{𝔹})  by composition with Ψ.

4.35. Theorem. Let A be a category. The composite A → 𝕊panA → ℙath_{∗}𝕊panA induces, by composition, an equivalence of categories,

GregDoub(ℙath_{∗}𝕊panA, 𝔹) ⟶^{≃} Cat(A, Vrt (𝔹)),

for any gregarious pseudo double category 𝔹, i.e., ℙath_{∗}𝕊pan: Cat → GregDoub is a left biadjoint of Vrt : GregDoub → Cat.

Proof. By the previous theorem,

psDoub(ℙath_{∗}𝕊panA, 𝔹) ⟶^{≃} OplaxN(𝕊panA, 𝔹).

The oplax double category 𝕊panA is gregarious and normal morphisms between gregarious double categories are automatically gregarious, so

OplaxN(𝕊panA, 𝔹) = GregOplax(𝕊panA, 𝔹) ⟶^{≃} Cat(A, Vrt (𝔹)),

by Theorem 4.25.

We denote the composition of functors ℙath_{∗}𝕊pan by Π_{2}. It is the double category version of the Π_{2}-construction studied in [14]. When we take the bicategory of horizontal arrows and special cells in ℙath_{∗}𝕊panA we obtain a bicategory which is biequivalent to Π_{2}A as introduced before.

Acknowledgement. We would like to thank the anonymous referee for various helpful comments.

References

[1] A. Bastiani, C. Ehresmann, Multiple functors. I. Limits relative to double categories, Cahiers Topologie Géom. Différentielle Catég. 15 (1974), pp. 215–292

[2] J. Bénabou, Introduction to bicategories, in Reports of the Midwest Category Seminar, Springer L.N.M. 47 (1967), pp. 1–77.

[3] R. Blackwell, G.M. Kelly, A.J. Power, Two-dimensional monad theory, J. Pure Appl. Algebra 59 (1989), pp. 1–41.

[4] Ronald Brown, Ghafar Mosa, Double categories, 2-categories, thin structures and connections, Theory and Applications of Categories 5 (1999), pp. 163–175.

[5] Albert Burroni, T -catégories (catégories dans un triple), Cahiers Topologie Géom. Différentielle Catég. 12 (1971), pp. 215–321.

[6] A. Carboni, G.M. Kelly, R.J. Wood, A 2-categorical approach to change of base and geometric morphisms I, Cahiers Topologie Géom. Différentielle Catég. 32 (1991), pp. 47–95.

[7] A. Carboni, G.M. Kelly, D. Verity, R.J. Wood, A 2-categorical approach to change of base and geometric morphisms, II, Theory and Applications of Categories 4 (1998), pp. 82–136.

[8] M.M. Clementino, D. Hofmann, W. Tholen, Exponentiability in categories of lax algebras, Theory and Applications of Categories 11 (2003), pp. 337–52.

[9] B.J. Day, P. McCrudden, R. Street, Dualizations and antipodes, Appl. Categ. Struct. 11 (2003), pp. 229–260.

[10] W.G. Dwyer, D.M. Kan, Calculating simplicial localizations, J. Pure Appl. Algebra 18 (1980), pp. 17–35.

[11] Robert Dawson, Robert Paré, General associativity and general composition for double categories, Cahiers Topologie Géom. Différentielle Catég. 34 (1993), no. 1, pp. 57– [12] R.J.MacG. Dawson, R. Paré, D.A. Pronk, Paths in double categories, Theory and

Applications of Categories 19 (2006), pp. 460–521.

[13] R.J.MacG. Dawson, R. Paré, D.A. Pronk, Universal properties of Span, in The Carboni Festschrift, Theory and Applications of Categories 13 (2005), pp. 61–85.

[14] Robert Dawson, Robert Paré, Dorette Pronk, Adjoining adjoints, Adv. Math. 178 (2003), pp. 99–140.

[15] R.J.MacG. Dawson, R. Paré, D.A. Pronk, Undecidability of the free adjoint construction, Applied Categorical Structures 11 (2003), pp. 403–419.

[16] Thomas M. Fiore, Simona Paoli and Dorette A. Pronk, Model structures on the category of small double categories, Algebraic and Geometric Topology 8 (2008), pp. 1855–1959.

[17] P. Gabriel, M. Zisman, Calculus of fractions and homotopy theory, Ergebnisse der Mathematik und ihrer Grenzgebiete, Band 35, Springer, New York (1967).

[18] Marco Grandis, Robert Paré, Adjoint for double categories, Cahiers Topologie Géom. Différentielle Catég. 45 (2004), no. 3, pp. 193–240.

[19] M. Grandis, R. Paré, Limits in double categories, Cahiers Topologie Géom. Différentielle Catég. 40 (1999), no. 3, pp. 162–220.

[20] John W. Gray, Formal Category Theory: Adjointness for 2-Categories, L.N.M. 391, Springer-Verlag, New York, 1974.

[21] C. Hermida, Representable multicategories, Adv. in Math. 151 (2000), pp. 164–225.

[22] C. Hermida, From coherent structures to universal properties, J. Pure Appl. Algebra 165 (2001), pp. 7–61.

[23] G.M. Kelly, R. Street, Review of the elements of 2-categories, in: Category Seminar, Springer L.N.M. 420, pp. 75–103.

[24] Stephen Lack, Icons, Applied Categorical Structures 18 (2010), pp. 289–307.

[25] J. Lambek, Deductive systems and categories II: Standard constructions and closed categories inCategory Theory, Homology Theory and their Applications, I (Battelle Institute Conference, Seattle, Wash., 1968, Vol. One), ed. P. Hilton, Springer L.N.M. 86 (1969), pp. 76–122.

[26] T. Leinster, Higher Operads, Higher Categories, London Mathematical Society Lecture Note Series 298, Cambridge University Press, 2003.

[27] Saunders Mac Lane, Robert Paré, Coherence for bicategories and indexed categories, J. Pure Appl. Algebra 37 (1985), no. 1, pp. 59–80.

[28] P. Palmquist, The double category of adjoint squares, Reports of the Midwest Category Seminar, V (Zürich, 1970), Springer L.N.M. 195 (1971), pp. 123–153.

[29] Jacques Penon, T -catégories représentables, Theory and Applications of Categories 22 (2009), pp. 376–387.

[30] M. Shulman, Framed bicategories and monoidal fibrations, Theory and Applications of Categories 20 (2008), pp. 650–738.

Department of Mathematics and Computing Science, Saint Mary’s University, Halifax, NS B3H 3C3, Canada and Department of Mathematics and Statistics, Dalhousie University, Halifax, NS B3H 3J5, Canada

Email: Robert.Dawson@smu.ca
      pare@mathstat.dal.ca
      pronk@mathstat.dal.ca

This article may be accessed at http://www.tac.mta.ca/tac/ or by anonymous ftp at

ftp://ftp.tac.mta.ca/pub/tac/html/volumes/24/13/24-13.{dvi,ps,pdf}

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods. Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication. Full text of the journal is freely available in .dvi, Postscript and PDF from the journal’s server at http://www.tac.mta.ca/tac/ and by ftp. It is archived electronically and in printed paper format.

Subscription information. Individual subscribers receive abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh, rrosebrugh@mta.ca.

Information for authors. The typesetting language of the journal is TEX, and L^{A}TEX2e strongly encouraged. Articles should be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files at http://www.tac.mta.ca/tac/.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: barr@math.mcgill.ca

Assistant TEX editor. Gavin Seal, Ecole Polytechnique Fédérale de Lausanne:

gavin seal@fastmail.fm

Transmitting editors.

Clemens Berger, Université de Nice-Sophia Antipolis, cberger@math.unice.fr Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown (at) btinternet.com Aurelio Carboni, Università dell Insubria: aurelio.carboni@uninsubria.it Valeria de Paiva, Cuill Inc.: valeria@cuill.com Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk P. T. Johnstone, University of Cambridge: ptj@dpmms.cam.ac.uk Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, Macquarie University: steve.lack@mq.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@acsu.buffalo.edu Tom Leinster, University of Glasgow, T.Leinster@maths.gla.ac.uk Jean-Louis Loday, Université de Strasbourg: loday@math.u-strasbg.fr Ieke Moerdijk, University of Utrecht: moerdijk@math.uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Brooke Shipley, University of Illinois at Chicago: bshipley@math.uic.edu James Stasheff, University of North Carolina: jds@math.unc.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Insubria: robert.walters@uninsubria.it R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca